### v1

# Kinetically Consistent Coarse Graining using Kernel-based Extended Dynamic Mode Decomposition

Abstract
In this paper, we show how kernel-based approximation to the Koopman generator – the kgEDMD algorithm – can be used to identify implied timescales and meta stable sets in stochastic dynamical systems, and to learn a coarse-grained dynamics on reduced variables, which retains the essential kinetic properties of the full model. The centerpiece of this study is a learning method to identify an effective diffusion in coarse-grained space by leveraging the kgEMD model for the Koopman generator. By combining this method with force matching, a complete model for the effective dynamics can be inferred. Using a two-dimensional model system and molecular dynamics simulation data of alanine dipeptide, we demonstrate that the proposed method successfully and robustly recovers the essential theormodynamic and kinetic properties of the full model. The parameters of the method can be determined using standard model validation techniques.
1 Introduction
Stochastic simulations of large-scale dynamical systems are widely used to model the behaviour of complex systems, with applications in computational physics, chemistry, materials science, and engineering. Many examples of such systems are high dimensional and subject to meta-stability, which means the system remains trapped in a set of geometrically similar configurations, while transitions to another such state are extremely rare. As a consequence, it becomes necessary to produce very long simulations in order to make statistically robust predictions. A prime example are atomistic molecular dynamics simulations (MD) 1 of macro-molecules, where meta-stability is typically caused by high energetic barriers separating deep potential energy minima 2 . As a result, it requires specialized high-performance computing facilities to reach the required simulation times, or it may just not be feasible at all 3, 4 .
Coarse graining (CG) describes the process of replacing the original dynamical system by a surrogate model on a (much) lower-dimensional space of descriptors 5, 6 , in such a way that certain properties of the original dynamics are preserved by the surrogate model. CG models can enable scientists to achieve much longer simulation times because of the reduced computational cost, while maintaining predictive capabilities of the full-order model. Setting up a CG model typically requires the following steps: first, the choice of a linear or non-linear mapping (CG map) from full state space to a lower-dimensional space, where the latter serves as the state space of the surrogate model. Second, definition of a parametric model class for the surrogate dynamics. Finally, determination of the parameters for that model class.
The first step is crucial to the CG model’s success, and has been a very active area of research for a long time, see Refs. 7, 8, 4 for reviews on this topic. In this study, we are concerned with the second and mainly the third step. CG models have often been parameterized using physically intuitive functional forms of, for instance, the coarse-grained energy. More recently, much more general functional forms have been used for the CG parameters, which are then approximated by powerful model classes, such as deep neural networks or reproducing kernels 9, 10, 11 , which is the approach we follow in this paper. We focus on CG for reversible stochastic differential equations (SDE) with a Boltzmann-type invariant distribution, which includes most popular simulation engines in MD, e.g. Langevin dynamics. Following the projection approach from Refs. 12, 13 , we likewise parameterize the coarse-grained model as a reversible SDE, disregarding memory terms.
Concerning parameter selection, the success of machine learning (ML) in recent years has led to the development of many powerful learning schemes for the parameters of a CG model. Examples are free energy learning 14 , and force matching 15 , among others. Many of these learning methods are geared towards ensuring thermodynamic consistency, which means that the surrogate model is trained to sample the marginalized Boltzmann distribution in CG space, thus ensuring accurate estimation of average quantities. Ensuring faithful reproduction of kinetic properties, such as time-correlation functions or transition time scales, is a much less developed topic, see e.g. 16, 17 for prior works in this area. In this paper, we focus on recovery of the transition time scales associated to meta stable states.
Transition rates and time scales are directly related to the leading spectrum of the system’s transfer or Koopman operator 18, 19, 20, 21 , which is the linear operator propagating expectation values of observable functions in time. Equivalently, one may also consider the spectrum of the associated Koopman generator (Kolmogorov operator for SDEs) close to zero. This connection has been at the heart of the Markov state modeling (MSM) approach 22, 23, 24 and many important developments based on it 25, 26, 27 . The spectral matching approach 28 makes use of this connection, by first parameterizing the CG model as a linear expansion using a fixed set of basis functions, and then solving a regression problem to recover the eigenvalues of the Koopman generator. This idea was formalized in Ref. 29 , by suggesting to regress on a full matrix representation of the Koopman generator, to capture as much kinetic content as possible. The generator matrix can be estimated a priori by a data-driven algorithm called generator EDMD (gEDMD).
In this study, we significantly improve the generator-based spectral matching, by formulating the approach using kernel methods. Reproducing kernels 30, 31 offer powerful model classes which have been applied in many areas of machine learning. Kernel-based approximation of the Koopman generator (kgEDMD) was first described in Ref. 32 We use the kgEDMD matrix and a (reduced and whitened) basis of kernel functions to formulate a learning problem for a generally state-dependent effective diffusion, which optimally approximates the Koopman generator. Our approach only requires simulation data of the full system, the choice of the kernel function, and measurements of the local diffusion, which is analogue to the local mean force in force matching. Combined with the latter, this approach provides complete access to the effective dynamics associated to a reversible system.
The key contributions of this work are:
  • We formulate a data-based learning problem for the effective diffusion of a coarse-grained SDE. The only hyper-parameters to be tuned are those of the kernel function. The method is robust to statistical noise and ill-conditioning as it is based on a whitened and truncated kernel bases.
  • We show that diffusion learning and force matching can be combined into a single pipeline to completely parameterize a coarse-grained SDE for reversible systems.
  • We show that thermodynamic and kinetic consistency are achieved by the method using two test cases, a two-dimensional model system and molecular dynamics simulations of the alanine dipeptide.
The structure of the paper is as follows: we introduce the required background on SDEs, coarse graining, and Koopman operator learning in Section 2. Our kernel-based learning framework is then presented in Section 3, while the numerical examples follow in Section 4. An overview of the notation is provided in Table 1, supplementary information on simulation details and model selection is given in the Appendix.
2 Theory
In this section, we provide the necessary background on stochastic dynamics, data-driven modeling, and Koopman spectral theory.
2.1 Stochastic processes
We consider a dynamical system described by a stochastic differential equation (SDE)
$dX_{t}=b(X_{t})dt+\sigma(X_{t})dW_{t},$ (1)
where $b(X_{t}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift vector field, $\mathbf{\sigma}(X_{t}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is the diffusion field, and $W_{t}$ is a d-dimensional Brownian motion. We sometimes refer to the diffusion covariance matrix which is denoted as $a\in\mathbb{R}^{d\times d}$:
$a(x)=\sigma^{T}(x)\sigma(x).$ (2)
Processes like Equation (1) are referred to as diffusion processes. A standard example commonly used in molecular modeling is overdamped Langevin dynamics
$dX_{t}=-\dfrac{1}{\gamma}\nabla F(X_{t})dt+\sqrt{2\beta^{-1}\gamma^{-1}}dB_{t},$ (3)
where $F:\Omega\rightarrow\mathbb{R}$ is the potential energy, $\beta=(k_{B}T)^{-1}$ and $\gamma$ are constants corresponding to the inverse temperature and the friction, respectively. The invariant measure for $X_{t}$ in Equation (3) is the Boltzmann distribution $\mu\propto\exp(-\beta F)$, and the dynamics are reversible with respect to $\mu$. More generally, a reversible SDE with invariant measure $\mu\propto\exp(-F)$ can be parameterized in terms of the generalized scalar potential $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, and the diffusion covariance $a$, as follows 33
$dX_{t}=\left[-\dfrac{1}{2}a(X_{t})\nabla F(X_{t})+\dfrac{1}{2}\nabla\cdot a(X_{t})\right]dt+\sigma(X_{t})dW_{t}.$ (4)
We will mostly consider reversible SDEs in this paper, and make use of the parameterization in Equation (4) when formulating learning methods.
2.2 Koopman and its generator
Koopman theory 34, 35 lifts the dynamics in Equation (1) into an infinite-dimensional space of observable functions to express the dynamics linearly. More precisely, the family of Koopman operators ${\mathcal{K}^{t}}$ for stochastic dynamics is defined as
$\mathcal{K}^{t}\psi(x)=\mathbb{E}^{x}\left[\psi(X_{t})\right]=\mathbb{E}\left[\psi(X_{t})\,|\,X_{0}=x\right],$ (5)
where $\mathbb{E}\left[\cdot\right]$ denotes the expected value. The operators $\mathcal{K}^{t}$ form a semigroup, that is $\mathcal{K}^{s+t}=\mathcal{K}^{s}\mathcal{K}^{t}$, and one can define their infinitesimal generator $\mathcal{L}$ as the time-derivative of the expectation value
$\mathcal{L}\psi(x)=\frac{d}{dt}\mathbb{E}^{x}\left[\psi(x_{t})\right]|_{t=0}.$ (6)
Stochastic calculus shows that for smooth observable functions, the generator can be written as 33
$\begin{split}\mathcal{L}\psi(x)&=b(x)\cdot\nabla\psi(x)+\frac{1}{2}a(x):\nabla^{2}\psi(x)\\ &=\sum_{i=1}^{d}b_{i}(x)\frac{\partial}{\partial x_{i}}\psi(x)+\frac{1}{2}\sum_{i,j=1}^{d}a_{ij}(x)\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\psi(x),\end{split}$ (7)
where $a$ and $b$ are the diffusion and drift terms defined above, $\nabla^{2}[\cdot]$ is the Hessian matrix of a function, and the colon :::: is a short-hand for the dot product between two matrices, also called Frobenius inner product. For overdamped Langevin dynamics, Eq. (7) simplifies to
$\mathcal{L}\psi(x)=-\frac{1}{\gamma}\nabla F(x)\cdot\nabla\psi(x)+\frac{1}{\gamma\beta}\Delta\psi(x).$
2.3 Spectral decomposition
A key quantity of interest are the eigenvalues and eigenfunctions of the generator. The study of spectral components of the generator helps us identify the long-time dynamics of the system. In molecular dynamics, we expect to find a number of eigenvalues close to zero, followed by a spectral gap. These low-lying eigenvalues are indicating the number of meta stable states of the system, which are the macrostates the system stays in the longest 18 . We write the eigenvalue problem for the generator as
$-\mathcal{L}\psi_{i}=\lambda_{i}\psi_{i}.$ (8)
The eigenvalues $\lambda_{i}$ of $-\mathcal{L}$ must be non-negative, and the lowest eigenvalue $\lambda_{1}=0$ is non-degenerate under very mild conditions 36 $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq...$ We also refer to the eigenvalues as rates, and to their reciprocals as implied timescales 22
$t_{i}=\frac{1}{\lambda_{i}}.$ (9)
2.4 Coarse graining
One of the main motivations of this work is to learn an SDE representing the full dynamics (1) on a coarse-grained space. Coarse graining (CG) is realized by mapping the state space $\Omega$ onto a lower-dimensional space $\hat{\Omega}\subset\mathbb{R}^{d_{\xi}}$ by means of a smooth CG function $\xi$. We write $\nu$ for the marginal distribution of the full-space invariant measure $\mu$ projected onto CG space. An important object is the space $L^{2}_{\nu}$ of functions on $\hat{\Omega}$ which are square integrable with respect to the weight function $\nu$. The space $L^{2}_{\nu}$ is a subspace of $L^{2}_{\mu}$ comprising all functions that depend only on the value of $\xi$. Moreover, the orthogonal projector onto $L^{2}_{\nu}$ from $L^{2}_{\mu}$ is the conditional expectation operator 12, 13
$\mathcal{P}\psi(z)=\frac{1}{\nu(z)}\mathbb{E}^{\mu}[\psi(x)|\xi(x)=z],$ (10)
where $z$ is a position in CG space. This operator calculates the average of a function $\psi$ over all $x\in\Omega$ whose projection onto CG space is the same point $z\in\hat{\Omega}$.
Considering $\mathcal{P}$, following the exposition in 13, one can define the projected generator
$\mathcal{L}^{\xi}=\mathcal{P}\mathcal{L}\mathcal{P}.$ (11)
It turns out its action on a function $\varphi\in L^{2}_{\nu}$ is given by
$\mathcal{L}^{\xi}(\phi)=\mathcal{P}[\mathcal{L}\xi]\cdot\nabla_{z}\phi+\frac{1}{2}\mathcal{P}[\nabla\xi^{T}a\nabla\xi]:\nabla_{z}^{2}\phi.$ (12)
As one can see, $\mathcal{L}^{\xi}$ is of the same form as the original generator $\mathcal{L}$ in Equation (7), and indeed it is the generator of an SDE $Z_{t}$ on $\hat{\Omega}$
$dZ_{t}=b^{\xi}(Z_{t})dt+\sigma^{\xi}(Z_{t})dW_{t}.$ (13)
The effective drift and diffusion coefficients are given in analytical form by
$b^{\xi}(z)=\mathcal{P}\left(\mathcal{L}\xi\right)(z)\hskip 28.45274pta^{\xi}(z)=\mathcal{P}\left(\nabla\xi^{T}a\nabla\xi\right)(z),$ (14)
and the practical task of coarse graining is to approximate them numerically. An important result, shown in 13 , and that we will use later on, is that the inner product of the projected generator with respect to the invariant measure $\nu$ in CG space is equal to the inner product of the original generator with respect to the invariant measure $\mu$ of the full space
$\left\langle\mathcal{L}\psi\circ\xi,\,\varphi\circ\xi\right\rangle_{\mu}=\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}$ (15)
where $\psi$ and $\varphi$ are functions in $L^{2}_{\nu}$.
2.5 Dynamic mode decomposition and its extension
Dynamic Mode Decomposition (DMD) is a conceptually simple method that provides a linear map to propagate the state of a dynamical system in time. The method was originally proposed in Ref. 37 and since then, many variations and extensions have been proposed, see Ref. 38, 39 for an overview. Basically, DMD looks for the best-fit linear operator $\mathbf{A}$ that advances the system state by solving the minimization problem in Equation (16)
$\mathbf{A}=\operatorname*{arg\,min}_{\mathbf{A}^{\prime}}\|\mathbf{X}^{\tau}-\mathbf{A}\mathbf{X}\|_{F}=\mathbf{X}^{\tau}\mathbf{X}^{\dagger}$ (16)
where $\mathbf{X}$ and $\mathbf{X}^{\tau}$ are collections of time-shifted snapshots of the dynamics, and $\dagger$ denotes the pseudo-inverse. Although DMD is a popular and widely-used method in data-driven modeling, it can only provide a linear model and the systems we are interested in for molecular dynamics are highly nonlinear.
Improved results can be obtained by using nonlinear functions instead of just the raw state variables. This method is known as Extended DMD (EDMD) 40 , that can provide more accurate models by capturing nonlinearities. After defining a dictionary of observables, $\{\psi_{i}\}_{i=1}^{n}$ (that are functions of the state space $\Omega$), we compile the evaluations of all dictionary functions at all snapshots into matrices
$\mathbf{\Psi}$ $=\left[\psi_{i}(x_{l})\right]_{i,l}\in\mathbb{R}^{n\times m},$ $\mathbf{\Psi}^{\tau}$ $=\left[\psi_{i}(x_{l+1})\right]_{i,l}\in\mathbb{R}^{n\times m},$
where we assume for simplicity that subsequent snapshots are separated by time $\tau$. We then solve the optimization problem:
$\mathbf{K}=\operatorname*{arg\,min}_{\mathbf{K^{\prime}}}\|\mathbf{\Psi}^{\tau}-\mathbf{K}^{\prime}\mathbf{\Psi}\|_{F}=\mathbf{\Psi}^{\tau}\mathbf{\Psi}^{\dagger}.$ (17)
The solution of Equation (17) can be obtained numerically using singular value decomposition. In the limit of infinite data, EDMD constitutes a finite-dimensional projection (Galerkin projection) of the Koopman operator onto the subspace spanned by the dictionary functions 40 .
2.6 Generator EDMD
The infinitesimal generator of the Koopman operator can also be approximated by an algorithm similar to EDMD, called gEDMD 29 . Once again, given a set of scalar basis functions $\psi(x)=\{\psi_{1}(x),...,\psi_{n}(x)\}$, and training data $\{x_{l}\}_{l=1}^{m}$, we form the matrices
$\mathbf{\Psi}$ $=\left[\psi_{i}(x_{l})\right]_{i,l},$ $\mathcal{L}\mathbf{\Psi}$ $=\left[\mathcal{L}\psi_{i}(x_{l})\right]_{i,l},$
using the analytical formula (7) to evaluate the second of these matrices. Solving an analogous regression problem to (17) leads to the matrix
$\mathbf{L}^{\top}=(\mathcal{L}\mathbf{\Psi})\mathbf{\Psi}^{\dagger}=\hat{\mathbf{A}}\hat{\mathbf{G}}^{-1},$ (18)
where
$\begin{gathered}\hat{\mathbf{A}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}(\mathcal{L}\psi_{i})(x_{k})\psi_{j}(x_{k}),\\ \hat{\mathbf{G}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\psi_{j}(x_{k}).\end{gathered}$ (19)
As for EDMD, it was proved in Ref. 29 that in the infinite-data limit, gEDMD converges to the Galerkin projection of the generator onto the space spanned by the basis functions
$\mathbf{L}^{\top}$ $=\mathbf{A}\mathbf{G}^{-1},$ $\mathbf{A}_{ij}$ $=\left\langle\mathcal{L}\psi_{i},\,\psi_{j}\right\rangle_{\mu}$ $\mathbf{G}_{ij}$ $=\left\langle\psi_{i},\,\psi_{j}\right\rangle_{\mu}.$ (20)
The matrices $\mathbf{G}$ and $\mathbf{A}$ are often called mass and stiffness matrix, respectively. For arbitrary stochastic dynamics, the computation of $\mathbf{A}$ involves a second-order differentiation as shown in Equation (7). However, if the stochastic dynamics are reversible, only first-order derivatives are required to compute the matrix $\mathbf{A}$, as the generator satisfies the following integration-by-parts formula
$\mathbf{A}_{ij}=\left<\mathcal{L}\psi_{i},\psi_{j}\right>_{\mu}=-\frac{1}{2}\int\nabla\psi_{i}\sigma\sigma^{T}\nabla\psi^{T}_{j}\,\mathrm{d}\mu.$ (21)
If the basis functions are actually defined in a CG space $\hat{\Omega}$, that is $\psi_{i}(x)=\psi_{i}(\xi(x))$, then by the chain rule the matrix $A$ can be written as
$\begin{split}\mathbf{A}_{ij}=-\frac{1}{2}\int\nabla_{x}\psi_{i}\sigma\sigma^{T}\nabla_{x}\psi^{T}_{j}\,\mathrm{d}\mu=-\frac{1}{2}\int(\nabla_{z}\psi_{i}\nabla_{x}\xi)\sigma\sigma^{T}(\nabla_{x}\xi^{T}\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu\\ =-\frac{1}{2}\int(\nabla_{z}\psi_{i})(\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T})(\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu.\end{split}$ (22)
We refer to the matrix
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ (23)
as local diffusion, and note that it is independent of the basis functions. It can therefore be computed a priori in numerical calculations.
3 Methods
3.1 Overview and Notation
Successful application of the gEDMD method requires selecting proper and expressive basis functions. This is not a trivial task and can highly depend on the dynamics of interest. To alleviate this, use of kernel methods has been proposed in Ref. 32 . The kgEDMD method is based on choosing the basis functions as features of a reproducing kernel $k(x,y)$ at the training data:
$\Phi=\left[k(\cdot,x_{1}),k(\cdot,x_{2}),...,k(\cdot,x_{m})\right]^{\top}.$ (24)
We consider $\Phi$ as a vector-valued function, that is, we write
$\Phi(x)=\left[k(x,x_{1}),k(x,x_{2}),...,k(x,x_{m})\right]^{\top}$ (25)
for the $m$-dimensional vector obtained by evaluating all the kernel features at $x$. Since kernels are two-argument functions, we can take their derivatives with respect to both arguments, and in fact we will need to distinguish these derivatives when formulating our methods. In general, we simply write $\nabla_{1}k$ and $\nabla_{2}k$ for gradients of the kernel function with respect to its first and second argument. Similarly, we write $\nabla_{1}\Phi,\,\nabla_{2}\Phi$ for the matrix-valued functions we obtain by assembling the gradients of all kernel features:
$\nabla_{1}\Phi(x)$ $=\begin{bmatrix}\nabla_{1}k(x,x_{1})\\ \vdots\\ \nabla_{1}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d},$ $\nabla_{2}\Phi(x)$ $=\begin{bmatrix}\nabla_{2}k(x,x_{1})\\ \vdots\\ \nabla_{2}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d}.$ (26)
One of the widely used kernel functions is the Gaussian kernel given by
$k(x_{i},x_{j})=\exp\left(-\frac{\|x_{i}-x_{j}\|)^{2}}{2\sigma^{2}}\right),$ (27)
where $\sigma$ is the standard deviation or bandwidth of the Gaussian kernel. If the state space is a periodic domain, we make use of a periodic Gaussian kernel instead 41
$k(x_{i},x_{j})=\exp\left(-\frac{\sum_{i=1}^{d}\sin\left(\frac{\pi}{p}(x_{i}-x_{j})\right)^{2}}{2\sigma^{2}}\right),$ (28)
where $\sigma$ is again the bandwidth of the Gaussian kernel, $p$ is the half-period of the domain, and $d$ is the dimension of system.
3.2 Kernel-based Koopman generator EDMD
Using the kernel feature basis as in Equation (24), we can seek a data-driven approximation to the eigenfunctions of the Koopman generator in the form $\psi_{i}(x)=\mathbf{u}_{i}^{\top}\Phi(x)$. As detailed in Ref. 32 , the coefficient vectors $\mathbf{u}_{i}$, are then found as solutions of the following generalized matrix eigenvalue problem
$\hat{\mathbf{A}}\mathbf{u}_{i}:=\frac{1}{2}\sum_{l=1}^{d}(\widehat{\mathbf{A}}^{(l)})^{\top}\widehat{\mathbf{A}}^{(l)}\mathbf{u}_{i}=\widehat{\lambda}_{i}\widehat{\mathbf{G}}\widehat{\mathbf{G}}\mathbf{u}_{i},$ (29)
where $d$ is the dimension of the state space. The matrices $\hat{\mathbf{A}}^{(l)}$ and $\hat{\mathbf{G}}$ are based on the reversible formulation for gEDMD, and essentially contain pairwise evaluations of the kernel and its first-order derivatives
$\begin{gathered}\left[\widehat{\mathbf{G}}\right]_{ns}=k(x_{n},x_{s})\\ \left[\widehat{\mathbf{A}}^{(l)}\right]_{ns}=\sigma_{l}(x_{n})^{\top}\nabla_{2}k(x_{n},x_{s}).\end{gathered}$ (30)
The kernelized mass matrix $\hat{\mathbf{G}}$ is often highly ill-conditioned. Therefore, we compute a reduced generator by performing a whitening transformation based on removing small eigenvalues of the mass matrix
$\widehat{\mathbf{G}}$ $=\mathbf{U}\Sigma\mathbf{U}^{T},$ $\mathbf{R}$ $=\mathbf{U}\Sigma^{-0.5}\in\mathbb{R}^{m\times r},$ $\hat{\mathbf{L}}_{r}$ $=\mathbf{R}^{T}\hat{\mathbf{A}}\hskip 2.84544pt\mathbf{R},$ (31)
in which $r\leq m$. Here, $\mathbf{R}$ is a transformation matrix mapping the original basis built by the feature map $\Phi$, to the reduced basis
$\mathbf{h}(x)=\mathbf{R}^{\top}\Phi(x)\in\mathbb{R}^{r}.$ (32)
In preparation for later, we also introduce the notation
$\nabla\mathbf{h}(x)=\left[\partial_{j}h_{l}(x)\right]_{l,j}=\mathbf{R}^{\top}\nabla_{1}\Phi(x)\in\mathbb{R}^{r\times d}$ (33)
for the gradients of all reduced features evaluated at point $x$. Diagonalization of the reduced generator matrix $\hat{\mathbf{L}}_{r}$ then provides the final estimates for the generator’s eigenvalues and eigenfunctions.
3.3 System identification
The reduced generator matrix $\hat{\mathbf{L}}_{r}$ provides an empirical model for the Koopman generator $\mathcal{L}$ on the space of whitened and rank-reduced features $\mathbf{h}$. Hence we can use this model towards other ends besides identification of the meta stable sets. Among other purposes, we can use it to identify an effective SDE governing the CG dynamics as in Equation (13), using a procedure that was already sketched in Ref. 29 . We parameterize the CG dynamics as a reversible SDE, which is justified if the original dynamics are reversible 13 . As discussed in Section 2 and in Ref. 33 , reversible systems can be characterized by the diffusion $a^{\xi}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ and a scalar potential $F^{\xi}:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}$, from which the drift is then obtained by:
$b^{\xi}=-\dfrac{1}{2}a^{\xi}\nabla_{z}F^{\xi}+\dfrac{1}{2}\nabla\cdot a^{\xi}.$ (34)
Thus, to identify the SDE, we need to learn these two components, which can be broken up into two separate problems. From now on, we employ a kernel $k$ on the CG space $\hat{\Omega}$, that is $k=k(z,z^{\prime})$.
3.3.1 Force matching
Knowing the full-process scalar potential field $F$ and the CG function $\xi$, the effective force can be obtained by a well-known technique called force matching 15 via the following minimization problem
$\nabla_{z}F^{\xi}=\underset{g\in L^{2}_{\nu}}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|g(\xi(x_{i}))-f^{\xi}_{\text{lmf}}(x_{i})\right\|^{2},$ (35)
where $m$ is the data size and $f^{\xi}_{\text{lmf}}$ is called local mean force and defined as follows
$\begin{gathered}f^{\xi}_{\text{lmf}}=-\nabla_{x}F\cdot G^{\xi}+\nabla_{x}\cdot G^{\xi}\\ G^{\xi}=\nabla_{x}\xi[(\nabla_{x}\xi)^{T}\nabla_{x}\xi]^{-1}.\\ \end{gathered}$ (36)
In line with the kernel-based approach we follow in this work, we parameterize the optimal vector field $g$ as the gradient of a linear combination of the feature basis $\Phi$ in Equation (24)
$F^{\xi}(z)$ $=\alpha^{\top}\Phi(z),$ $g(z)$ $=\alpha^{\top}\nabla_{1}\Phi(z)=\sum_{i=1}^{m}\alpha_{i}\nabla_{1}k(z,z_{i}).$ (37)
The force matching problem thus breaks down to determination of $m$ optimal expansions coefficients $\alpha$. We stress, however, that any other parameterization could be used for this step, for example a deep learning-based method 11 .
3.3.2 Kernel-based diffusion learning
Our main contribution is a learning method for the effective diffusion, complementing the effective potential $F^{\xi}$. We first use training data of the full space dynamics and the kernel $k$ on the CG space $\hat{\Omega}$, to learn a reduced matrix representation $\hat{\mathbf{L}}_{r}$ of the full generator. Recalling Equation (15), the matrix $\hat{\mathbf{L}}_{r}$ is at the same time an approximation to the CG generator $\mathcal{L}^{\xi}$. On the CG space, the reversible generator $\mathcal{L}^{\xi}$ also satisfies the integration-by-parts formula (21),
$\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}=-\frac{1}{2}\int_{\hat{\Omega}}\nabla_{z}\psi(z)a^{\xi}(z)\nabla_{z}\varphi(z)^{\top}\,\mathrm{d}\nu(z),$ (38)
which only features an explicit dependence on the unknown effective diffusion $a^{\xi}$, see Equation (14). Using the reduced basis $\mathbf{h}(z)=\mathbf{R}^{\top}\Phi(z)$, we parameterize the effective diffusion as
$\sigma^{\xi}_{\alpha}(z)=\sum^{r}_{l=1}\alpha_{l}h_{l}(z),$ (39)
where, in the most general case, $\alpha_{l}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ must be matrix-valued coefficients. To arrive at a compact expression, we gather all unknown expansion coefficients $\alpha_{l}$ as a third order tensor $\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}$, which allows us to write
$\sigma^{\xi}_{\alpha}(z)=\alpha\cdot_{|3,1}\mathbf{h}(z)=\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (40)
using the notation $\mathbf{a}\cdot_{|i,j}\mathbf{b}$ to denote contraction of two tensors $\mathbf{a},\,\mathbf{b}$ over indices $i$ and $j$ of commensurate dimension. Plugging this into Equation (38) and using the data-driven estimator of Equation (30), we obtain a parametric model for the reduced generator matrix $\hat{\mathbf{L}}_{r}$
$\begin{split}\left[\hat{\mathbf{L}}^{\xi}_{\alpha}\right]&=\dfrac{1}{2m}\sum_{k=1}^{m}\nabla\mathbf{h}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})^{\top}\nabla\mathbf{h}^{\top}(z_{k})\\ &=\dfrac{1}{2m}\sum_{k=1}^{m}\mathbf{S}^{\alpha}(z_{k})\mathbf{S}^{\alpha}(z_{k})^{\top},\\ \mathbf{S}^{\alpha}(z)&:=\nabla\mathbf{h}(z)\sigma_{\alpha}^{\xi}(z).\end{split}$ (41)
Combining Equations (33) and (40), we can compactly write
$\mathbf{S}^{\alpha}(z)=\mathbf{R}^{\top}\nabla_{1}\Phi(z)\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (42)
which is now expressed solely in terms of the kernel features $\Phi$, the whitening transformation $\mathbf{R}$, and the unknown coefficient tensor $\alpha$. We then learn the effective diffusion by matching the parameterized generator matrix $\hat{\mathbf{L}}^{\xi}_{\alpha}$ to the true reduced generator matrix $\hat{\mathbf{L}}_{r}$, therefore solving the following minimization problem:
$\alpha^{*}=\underset{\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}}{\arg\min}\left[\left\|\hat{\mathbf{L}}_{r}-\hat{\mathbf{L}}^{\xi}_{\alpha}\right\|_{2}^{2}+\gamma\left\|\frac{\partial\sigma}{\partial z}\right\|_{2}^{2}\right],$ (43)
where $\hat{\mathbf{L}}^{\xi}_{\alpha}$ is calculated from $\alpha$ via Equations (41) and (42). To avoid over-fitting and to suppress oscillations in $\sigma$, we found that introducing the second term to regularize the derivatives of the diffusion provides the best results.
Here, the effective diffusion is represented by a full matrix. However, for simplicity, we only learn a diagonal diffusion matrix for the numerical results reported in the following section. The method can be summarized in Algorithm 1.
Algorithm 1 Learning Effective Dynamics with kgEDMD
Input: full space data $\{x_{k}\}_{k=1}^{m}$ in $\mathbb{R}^{d}$, CG map $\xi$, kernel function $k$,
Truncation rule for $r$ in Equation (31), regularization $\gamma$
1:Diffusion learning
2:     Compute local diffusion $a^{\xi}_{\rm loc}$ as in Equation (23).
3:     Form mass and stiffness matrices $\mathbf{G}$ and $\mathbf{A}^{(l)}$ as in Equation (30)
4:     Compute reduced generator $\hat{\mathbf{L}}_{r}$ as in Equation (31)
5:     Perform the minimization problem as in Equation (43)
6:     Form the effective diffusion $\sigma^{\xi}_{\alpha}$ as in Equation (39)
7:Force matching
8:     Compute the local mean force $f^{\xi}_{\text{lmf}}$ as in Equation (36)
9:     Perform the minimization problem as in Equation (35)
10:     Form the effective potential $F^{\xi}$ as in Equation (37)
4 Examples
To show the effectiveness of the proposed method, we apply it to a two-dimensional model system defined by the Lemon-slice potential, and to MD simulation data of the alanine dipeptide, which is a widely-used test case in molecular dynamics.
As the optimization block for diffusion learning is non-convex, we use a library function from scipy’s optimization toolbox, but solve the problem repeatedly for $50$ randomly generated initial guesses. The optimization block for the force matching, on the other hand, is done via ridge regression.
4.1 Lemon-Slice potential
4.1.1 System introduction
The Lemon-slice system is governed by overdamped Langevin dynamics in Equation (3) with the following potential $F$
$F(x,y)=F(r,\phi)=\cos(4\phi)+10(r-1)^{2},$ (44)
where $r$ and $\phi$ are polar coordinates. The energy landscape of the system is shown in Figure 1 a. To form the SDE for this example, we consider a diagonal state-dependent diffusion field $\sigma(x)$ defined as
$\sigma(x)=\begin{bmatrix}\sqrt{\frac{2}{\beta}(\sin(\phi)+1.5)}&0\\ 0&\sqrt{\frac{2}{\beta}(\sin(\phi)+1.5)}\\ \end{bmatrix}$ (45)
where $\beta=1$ is the inverse temperature. Using the Euler-Maruyama scheme at discrete integration time step $dt=10^{-3}$ for integration of the SDE, we collect the training data for learning. For the sake of validation and showing the robustness of the method, we produce $5$ independent experiments, each with length of $m=10^{5}$ time steps. We further down sample them to $1000$ samples each for learning effective force and diffusion.
As shown in previous studies, the polar angle $\phi$ is a suitable CG coordinate for this system, as it resolves all four meta stable states
$\xi(x,y)=\phi.$ (46)
For this system, analytical expressions for the effective drift and diffusion along $\xi$ can be obtained by a slight modification of the results in 42 , and serve as reference values.
We use the kgEDMD algorithm on the reaction coordinate $\xi$, to identify the generator eigenvalues and meta stable states and, subsequently, to identify an effective dynamics along $\xi$ using Algorithm 1.
As a first try, we consider the Gaussian kernel for Galerkin approximation of the generator. Notice that the CG dynamics of Lemon-slice problem is periodic in the polar angle, however, a generic Gaussian kernel is not so. Thus, it cannot capture the periodicity of the system of interest. As an alternative, we also use the periodic Gaussian kernel as in Equation (28) and analyze both. The kernel bandwidth in either versions of Gaussian kernel is optimized using cross validation based on the VAMP-score 27 . Details on the VAMP-score analysis are reported in the appendix.
4.1.2 Meta-stability analysis
Figure 1 c shows the leading eigenvalues obtained from the kgEDMD model. As one notices, there are four dominant eigenvalues followed by a gap. These four eigenvalues are corresponding to the four minima in the potential field. Having determined the eigenvectors of the generator, we can perform robust Perron Cluster Cluster Analysis (PCCA+) 43 algorithm to assign to each sample point its membership to each meta stable state. Figure 1 b shows that the four potential minima are perfectly recovered in this way.
Figure 1: Approximation of generator for the Lemon slice system. Potential field in (a). Membership analysis in (b) using $1000$ samples. The dominant eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion, using Gaussian and periodic Gaussian kernels, in (c). The relative error of these eigenvalues compared to the reference is shown in (d).
A comparison of the leading eigenvalues of the reference model $\hat{\mathbf{L}}_{r}$ and the optimal matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$ is shown in Figure 1 c. Both choices of the kernel function lead to satisfactory results, the periodic kernel provides slightly higher accuracy in approximation of the generator eigenvalues. Note that the kernel bandwidth is tuned for each kernel function separately.
4.1.3 Identification of CG dynamics
The learned generator providing the eigenvalues reported above is built upon the effective diffusion shown in Figure 2 b, which is almost perfectly following the reference. Furthermore, we perform the force matching as well and obtain the effective force in the CG space shown in Figure 2 a. From the effective force and diffusion, the effective drift can be obtained according to Equation (34), which is also compared against the analytical expression in Figure 2 c, likewise showing very good agreement.
Having effective drift and diffusion, we are able to build the learned SDE governing the CG coordinate. We use the Euler-Maruyama scheme to integrate the learned and reference SDEs with integration time step of $dt=10^{-3}$. Figure 2 d shows two trajectories of the CG coordinate $\phi$ for both dynamics for $10^{4}$ time steps, using the same Brownian motion for both trajectories. The propagated learned system follows the reference closely, with both systems staying long times in each meta stable state, and rarely swapping in between those. The results demonstrate that the proposed method can approximate the full system’s meta stable sets well, and identify a suitable SDE for CG dynamics which is accurate even on the level of individual trajectories.
Figure 2: Application of Algorithm 1 to identify angular dynamics for the Lemon-slice system. Effective force in (a), effective diffusion in (b), effective drift in (c), and integration of an example trajectory, using both the reference and learned SDE in (d).
In this example, both the full dynamics and the learned effective dynamics possess state-dependent diffusion fields. We compare the properties of the learned CG model with variable diffusion to those of a CG dynamics with constant diffusion. For the latter, we set the effective diffusion constant to $a=\frac{2}{\beta}=2$. We propagate the corresponding SDEs for a sufficiently large span of time, and estimate a new generator EDMD model based on these simulations, this time using Random Fourier Features as observables 44 . Figure 3, shows the eigenvalues of the generator for these cases compared to the learned generator built upon the original dataset. The result shows that learning a state-dependent diffusion is necessary to recover the original system’s leading eigenvalues. To show the scalability of the method, we repeat the analysis on a real system in the following section.
Figure 3: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (SDD, orange) and with constant diffusion (CD, green). As a comparison, we show the eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ using the original dataset (blue triangles). Note that the first eigenvalue is omitted as it is zero.
4.2 Alanine dipeptide
4.2.1 System introduction
[…]
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi(x)\frac{2}{\beta\gamma}M^{-1}\nabla_{x}\xi^{\top}(x)$ (47)
where $M$ is the diagonal mass matrix of all atoms, $\beta$ is the inverse temperature, $\gamma$ is the friction, and $\nabla_{x}\xi(x)$ is the Jacobian of dihedral angles. This way of computing the local diffusion corresponds to a reversible overdamped Langevin process in full position space, meaning that the resulting CG model is an approximation to that same overdamped process. Fitting an overdamped model is simpler than modeling the full non-reversible underdamped Langevin process if a non-linear CG map is used. In addition, the underdamped process is expected to be close to an overdamped model in position space after a re-scaling of time. However, this also means that the eigenvalues and timescales of the overdamped model are expected to differ from the ones obtained by directly approximating the underdamped Langevin model, while the meta stable states will remain the same.
The free energy landscape of the system with respect to these two angles is shown in Figure 4.
Figure 4: Graphical representation of the alanine dipeptide molecule on the left, and the reference free energy profile in two-dimensional dihedral angle space on the right.
The free energy landscape shows four minima, two on the left side, and two in the central part. We apply the kgEDMD algorithm to find the meta stable sets, and then use Algorithm 1 to learn the effective force and a state-dependent effective diffusion field in the dihedral angle space. Because of the periodicity of the CG coordinates, $\phi$ and $\psi$, we adopt the periodic Gaussian kernel. Similar to the previous example, we tune the bandwidth in the kernel function using the VAMP-score.
4.2.2 Meta stability analysis
Figure 5 a shows the leading eigenvalues of the generator obtained from the kgEDMD model.
Figure 5: Approximation of generator for alanine dipeptide. The dominant eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error of these eigenvalues compared to the reference is shown on the right.
The figure indicates the four dominant eigenvalues which are corresponding to the four minima in the free energy landscape followed by a gap. In addition, we also show the eigenvalues of the generator $\mathbf{L}^{\xi}_{\alpha}$ corresponding to the optimal effective diffusion, which agree well with the reference. Errorbars in the figures are generated by analyzing $5$ independent subsampled sets of the original data set, each one comprising $2000$ samples.
4.2.3 Identification of CG dynamics
For this 2-dimensional CG dynamics, we can express the diffusion field in a $2\times 2$ full matrix. Here we assume though, that the learned diffusion is a diagonal matrix. Figure 6 shows the first and second diagonal terms of the learned diffusion field.
Figure 6: The first (a) and second (b) diagonal terms of the learned diffusion covariance matrix $a$, the reference free energy surface (c) and the free energy surface learned via kernel-based force matching (d).
For the diffusion learning process, we incorporate $2000$ samples of the available dataset.
The next step is to apply force matching algorithm to obtain the effective force. The reference and learned effective free energy surface are depicted in Figures 6c and  6d, respectively. It it noticeable that the learned free energy surface correctly captures all energetic minima, and approximates the energy barriers properly. We also note that $m=50,000$ samples were required to obtain the results in Figure 6 d with kernel-based force matching.
[…]
Figure 8: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (Prop.-SDD, green) and with constant diffusion (Prop.-CD, red). As a comparison, we show the eigenvalues of the optimal learned generator matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$, assuming either state-dependent diffusion (blue dots) or constant diffusion (orange triangles). Note that the first eigenvalue is omitted as it is zero.
[…]
5 Discussion
We presented the use of kernel-based approximation to the Koopman generator – kgEDMD – to determine meta stable sets and to learn the parameters of CG models for stochastic systems. For reversible systems, we presented a single learning pipeline to identify both the potential energy and the diffusion governing the effective SDE in CG space. By means of two examples, a two-dimensional model potential and the alanine dipeptide, we showed that the learned effective dynamics is able to reproduce both thermodynamic and kinetic quantities of the full dynamics with high accuracy.
In the molecular example, we found that the CG dynamics based on the fully state-dependent diffusion performed comparably to a constant-diffusion model using the average value of the learned diffusion. It should be noted, however, that determining this average value requires to learn the optimal state-dependent model in the first place. However, these observations can serve as a basis for a bottom-up construction of the effective diffusion, starting from a simple constant model, and increasing the complexity incrementally. Additional future research topics include improvements to the efficiency of the method, using, for example, sparse approximations using random Fourier features 45 . Moreover, one can also try to simultaneously optimize the CG mapping $\xi$ along with the parameters of the CG model, for instance by balancing the VAMP score versus the complexity of the CG model.
Data Availability Statement
Codes and data to reproduce the results and figures shown in this manuscript are available from the following public repository: 10.5281/zenodo.13808938.
6 Appendix
6.1 Notation
The most important notation used in the manuscript is summarized in Table 1.
Table 1: Overview of notation
$X_{t}$ stochastic process
$k,\Phi$ kernel and associated feature map
$\mathbb{H}$ reproducing kernel Hilbert space
$\mathcal{K}^{\tau}$ Koopman operator with lag time $\tau$
$\mathcal{L}$ generator of the Koopman operator
$\mathbf{h}$ reduced basis set from whitening transformation
$\hat{\mathbf{L}}$, $\hat{\mathbf{L}}_{r}$ generator matrix and reduced generator matrix
$\sigma^{\xi}_{\alpha}$ effective diffusion parameterized by $\alpha$
$\hat{\mathbf{L}}^{\xi}_{\alpha}$ effective generator matrix for diffusion with parameters $\alpha$
$F$, $F^{\xi}$ potential and effective potential
$f^{\xi}_{\rm lmf}$, $a^{\xi}_{\rm loc}$ local mean force and local diffusion
$\nabla_{1}f$ gradient of $f$ with respect to the first argument
$\nabla_{2}f$ gradient of $f$ with respect to the second argument
$A\cdot_{|a,b}B$ contraction of arguments $a$-th of $A$ and $b$-th of $B$
6.2 VAMP-score
We tune hyper-parameters of the proposed method based on VAMP variational principle proposed in 27, stating that for reversible systems, the $k$ dominant eigenvalues of the Koopman generator can be obtained by a minimization problem
$\sum_{i=1}^{k}\lambda_{i}=\min_{\phi_{0},...,\phi_{k}}\sum_{i=1}^{k}\left\langle\phi_{i},\,\mathcal{L}\phi_{i}\right\rangle_{\mu}$ (48)
where the $\phi_{i}$ are orthogonal functions. We use this variational principle to optimize the kernel bandwidth in the context of our proposed method. To do this robustly and avoid overfitting, we make use of standard cross validation scheme by introducing $40\%$ of dataset as test set. Figure 9 shows the result for the lemon slice example, using Gaussian and periodic Gaussian kernels.
Figure 9: VAMP-score analysis for the Lemon slice example using periodic Gaussian kernel on the left and Gaussian kernel on the right
We applied the same procedure for optimizing the bandwidth for the example of alanine dipeptide. The only difference is that due to the relative complexity of this example compared to the Lemon slice, and requiring to incorporate more data for the optimization, we use random Fourier features (RFF) to reduce the size of basis functions, leading to a low-rank approximation of the generator. Figure 10 shows the result for optimization of the bandwidth.
Figure 10: VAMP-score analysis for alanine dipeptide.
6.3 Simulation Settings for Alanine Dipeptide
[…]
Table 2: Experiment setup
Force Field AMBER99SB-ILDN
[…]
Integrator Langevin dynamics
[…]

### v2

@@
 Abstract
-In this paper, we show how kernel-based approximation to the Koopman generator – the kgEDMD algorithm – can be used to identify implied timescales and meta stable sets in stochastic dynamical systems, and to learn a coarse-grained dynamics on reduced variables, which retains the essential kinetic properties of the full model.
-The centerpiece of this study is a learning method to identify an effective diffusion in coarse-grained space by leveraging the kgEMD model for the Koopman generator.
-By combining this method with force matching, a complete model for the effective dynamics can be inferred.
-Using a two-dimensional model system and molecular dynamics simulation data of alanine dipeptide, we demonstrate that the proposed method successfully and robustly recovers the essential theormodynamic and kinetic properties of the full model.
+In this paper, we show how kernel-based models for the Koopman generator – the gEDMD method – can be used to identify coarse-grained dynamics on reduced variables, which retain the slowest transition timescales of the original dynamics.
+The centerpiece of this study is a learning method to identify an effective diffusion in coarse-grained space, which is similar in spirit to the force matching method.
+By leveraging the gEDMD model for the Koopman generator, the kinetic accuracy of the CG model can be evaluated.
+By combining this method with a suitable learning method for the effective free energy, such as force matching, a complete model for the effective dynamics can be inferred.
+Using a two-dimensional model system and molecular dynamics simulation data of alanine dipeptide and the Chignolin mini-protein, we demonstrate that the proposed method successfully and robustly recovers the essential kinetic and also thermodynamic properties of the full model.
 The parameters of the method can be determined using standard model validation techniques.
@@
 A prime example are atomistic molecular dynamics simulations (MD) 1 of macro-molecules, where meta-stability is typically caused by high energetic barriers separating deep potential energy minima 2 .
-As a result, it requires specialized high-performance computing facilities to reach the required simulation times, or it may just not be feasible at all 3, 4 .
-Coarse graining (CG) describes the process of replacing the original dynamical system by a surrogate model on a (much) lower-dimensional space of descriptors 5, 6 , in such a way that certain properties of the original dynamics are preserved by the surrogate model.
+As a result, it requires specialized high-performance computing facilities to reach the required simulation times, or it may just not be feasible at all 3 .
+Coarse graining (CG) describes the process of replacing the original dynamical system by a surrogate model on a (much) lower-dimensional space of descriptors 4, 5 , in such a way that certain properties of the original dynamics are preserved by the surrogate model.
 CG models can enable scientists to achieve much longer simulation times because of the reduced computational cost, while maintaining predictive capabilities of the full-order model.
@@
 The first step is crucial to the CG model’s success, and has been a very active area of research for a long time, see Refs.
-7, 8, 4 for reviews on this topic.
-In this study, we are concerned with the second and mainly the third step.
-CG models have often been parameterized using physically intuitive functional forms of, for instance, the coarse-grained energy.
+6, 7, 8 for reviews on this topic.
+In this study, we only show examples of low-dimensional CG coordinates that have already been validated, while the problem of learning high-dimensional and fully transferrable CG models is left for future studies.
+Our focus is instead on the second and mainly the third step.
+CG models have often been parameterized using physically intuitive functional forms for the coarse-grained energy.
 More recently, much more general functional forms have been used for the CG parameters, which are then approximated by powerful model classes, such as deep neural networks or reproducing kernels 9, 10, 11 , which is the approach we follow in this paper.
-We focus on CG for reversible stochastic differential equations (SDE) with a Boltzmann-type invariant distribution, which includes most popular simulation engines in MD, e.g.
+We study CG for reversible stochastic differential equations (SDE) with a Boltzmann-type invariant distribution, which is applicable to most popular simulation engines in MD, e.g.
 Langevin dynamics.
@@
 12, 13 , we likewise parameterize the coarse-grained model as a reversible SDE, disregarding memory terms.
-Concerning parameter selection, the success of machine learning (ML) in recent years has led to the development of many powerful learning schemes for the parameters of a CG model.
+The success of machine learning (ML) in recent years has led to the development of many powerful learning schemes for the parameters of a CG model.
 Examples are free energy learning 14 , and force matching 15 , among others.
 Many of these learning methods are geared towards ensuring thermodynamic consistency, which means that the surrogate model is trained to sample the marginalized Boltzmann distribution in CG space, thus ensuring accurate estimation of average quantities.
-Ensuring faithful reproduction of kinetic properties, such as time-correlation functions or transition time scales, is a much less developed topic, see e.g.
-16, 17 for prior works in this area.
-In this paper, we focus on recovery of the transition time scales associated to meta stable states.
-Transition rates and time scales are directly related to the leading spectrum of the system’s transfer or Koopman operator 18, 19, 20, 21 , which is the linear operator propagating expectation values of observable functions in time.
+Ensuring faithful reproduction of kinetic properties, such as time-correlation functions or transition timescales, is a much less developed topic, see Ref.
+16 for an overview of prior work in this area.
+In this paper, we focus on the recovery of implied transition timescales 17 associated to meta stable states.
+Implied transition timescales are defined in terms of the leading spectrum of the system’s transfer or Koopman operator 18, 19, 20, 21 .
 Equivalently, one may also consider the spectrum of the associated Koopman generator (Kolmogorov operator for SDEs) close to zero.
-This connection has been at the heart of the Markov state modeling (MSM) approach 22, 23, 24 and many important developments based on it 25, 26, 27 .
-The spectral matching approach 28 makes use of this connection, by first parameterizing the CG model as a linear expansion using a fixed set of basis functions, and then solving a regression problem to recover the eigenvalues of the Koopman generator.
+This connection has been at the heart of the Markov state modeling (MSM) approach 17, 22, 23 and many important developments based on it 24, 25, 26 .
+The spectral matching approach 27 was the first to make use of this connection, by first parameterizing the CG model as a linear expansion of fixed basis functions, and then solving a regression problem to recover the eigenvalues of the Koopman generator.
 This idea was formalized in Ref.
-29 , by suggesting to regress on a full matrix representation of the Koopman generator, to capture as much kinetic content as possible.
+28 , by suggesting to regress on a full matrix representation of the Koopman generator.
 The generator matrix can be estimated a priori by a data-driven algorithm called generator EDMD (gEDMD).
-In this study, we significantly improve the generator-based spectral matching, by formulating the approach using kernel methods.
-Reproducing kernels 30, 31 offer powerful model classes which have been applied in many areas of machine learning.
-Kernel-based approximation of the Koopman generator (kgEDMD) was first described in Ref.
-32 We use the kgEDMD matrix and a (reduced and whitened) basis of kernel functions to formulate a learning problem for a generally state-dependent effective diffusion, which optimally approximates the Koopman generator.
-Our approach only requires simulation data of the full system, the choice of the kernel function, and measurements of the local diffusion, which is analogue to the local mean force in force matching.
-Combined with the latter, this approach provides complete access to the effective dynamics associated to a reversible system.
-The key contributions of this work are:
+In this study, we significantly improve on the idea of leveraging the Koopman generator for the identification of coarse grained models. Our key contributions are the following:
   • We formulate a data-based learning problem for the effective diffusion of a coarse-grained SDE.
+This formulation is analogous to the force matching approach for the coarse-grained energy.
+Just as force matching relies on measurements of the local mean force, our approach rests on a similar quantity called local diffusion.
+  • We suggest to parametrize the diffusion by a basis of random Fourier features 29 , which form a widely-used approximation technique for reproducing kernels.
+Random features offer a compromise between representational power and computational efficiency.
 The only hyper-parameters to be tuned are those of the kernel function.
-The method is robust to statistical noise and ill-conditioning as it is based on a whitened and truncated kernel bases.
-  • We show that diffusion learning and force matching can be combined into a single pipeline to completely parameterize a coarse-grained SDE for reversible systems.
-  • We show that thermodynamic and kinetic consistency are achieved by the method using two test cases, a two-dimensional model system and molecular dynamics simulations of the alanine dipeptide.
+The method is robust to statistical noise and ill-conditioning as it is based on a whitened and truncated basis set.
+  • We show how the gEDMD method can be leveraged to evaluate the kinetic consistency of the learned CG model by comparing its eigenvalues to those of the reference gEDMD matrix.
+This assessment does not require simulations of the CG model.
+  • We show that kinetic and also theromodynamic consistency are achieved by the method using three test cases, a two-dimensional model system and molecular dynamics simulations of the alanine dipeptide and the Chignolin mini-protein.
+For the molecular systems, we consider CG models in low-dimensional reaction coordinate spaces, and we employ an overdamped assumption to simplify the learning process.
 The structure of the paper is as follows: we introduce the required background on SDEs, coarse graining, and Koopman operator learning in Section 2.
-Our kernel-based learning framework is then presented in Section 3, while the numerical examples follow in Section 4.
+Our learning framework is then presented in Section 3, while the numerical examples follow in Section 4.
 An overview of the notation is provided in Table 1, supplementary information on simulation details and model selection is given in the Appendix.
@@
 We sometimes refer to the diffusion covariance matrix which is denoted as $a\in\mathbb{R}^{d\times d}$:
-$a(x)=\sigma^{T}(x)\sigma(x).$ (2)
-Processes like Equation (1) are referred to as diffusion processes. A standard example commonly used in molecular modeling is overdamped Langevin dynamics
+$a(x)=\sigma(x)\sigma^{\top}(x).$ (2)
+A standard example for Eq. (1), commonly used in molecular modeling, is overdamped Langevin dynamics
 $dX_{t}=-\dfrac{1}{\gamma}\nabla F(X_{t})dt+\sqrt{2\beta^{-1}\gamma^{-1}}dB_{t},$ (3)
@@
 The invariant measure for $X_{t}$ in Equation (3) is the Boltzmann distribution $\mu\propto\exp(-\beta F)$, and the dynamics are reversible with respect to $\mu$.
-More generally, a reversible SDE with invariant measure $\mu\propto\exp(-F)$ can be parameterized in terms of the generalized scalar potential $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, and the diffusion covariance $a$, as follows 33
+More generally, a reversible SDE with invariant measure $\mu\propto\exp(-F)$ can be parameterized in terms of the generalized scalar potential $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, and the diffusion covariance $a$, as follows 30
 $dX_{t}=\left[-\dfrac{1}{2}a(X_{t})\nabla F(X_{t})+\dfrac{1}{2}\nabla\cdot a(X_{t})\right]dt+\sigma(X_{t})dW_{t}.$ (4)
-We will mostly consider reversible SDEs in this paper, and make use of the parameterization in Equation (4) when formulating learning methods.
-2.2 Koopman and its generator
-Koopman theory 34, 35 lifts the dynamics in Equation (1) into an infinite-dimensional space of observable functions to express the dynamics linearly.
+We will only consider reversible SDEs in this paper, and make use of the parametrization in Equation (4) when formulating learning methods.
+2.2 Koopman generator and spectral decomppsition
+Koopman theory 31, 32 lifts the dynamics in Equation (1) into an infinite-dimensional space of observable functions to express the dynamics linearly.
 More precisely, the family of Koopman operators ${\mathcal{K}^{t}}$ for stochastic dynamics is defined as
@@
 where $\mathbb{E}\left[\cdot\right]$ denotes the expected value.
-The operators $\mathcal{K}^{t}$ form a semigroup, that is $\mathcal{K}^{s+t}=\mathcal{K}^{s}\mathcal{K}^{t}$, and one can define their infinitesimal generator $\mathcal{L}$ as the time-derivative of the expectation value
-$\mathcal{L}\psi(x)=\frac{d}{dt}\mathbb{E}^{x}\left[\psi(x_{t})\right]|_{t=0}.$ (6)
-Stochastic calculus shows that for smooth observable functions, the generator can be written as 33
+The associated infinitesimal generator $\mathcal{L}$ is the time-derivative of the expectation value, which can be written as a linear differential operator:
 $\begin{split}\mathcal{L}\psi(x)&=b(x)\cdot\nabla\psi(x)+\frac{1}{2}a(x):\nabla^{2}\psi(x)\\ &=\sum_{i=1}^{d}b_{i}(x)\frac{\partial}{\partial x_{i}}\psi(x)+\frac{1}{2}\sum_{i,j=1}^{d}a_{ij}(x)\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\psi(x),\end{split}$ (6)
-where $a$ and $b$ are the diffusion and drift terms defined above, $\nabla^{2}[\cdot]$ is the Hessian matrix of a function, and the colon :::: is a short-hand for the dot product between two matrices, also called Frobenius inner product.
+where $a$ and $b$ are the diffusion and drift terms defined above, $\nabla^{2}[\cdot]$ is the Hessian matrix of a function, and the colon :::: is a short-hand for the dot product between two matrices.
 For overdamped Langevin dynamics, Eq.
@@
 $\mathcal{L}\psi(x)=-\frac{1}{\gamma}\nabla F(x)\cdot\nabla\psi(x)+\frac{1}{\gamma\beta}\Delta\psi(x).$
-2.3 Spectral decomposition
-A key quantity of interest are the eigenvalues and eigenfunctions of the generator.
+The key quantity of interest are the eigenvalues and eigenfunctions of the generator.
 The study of spectral components of the generator helps us identify the long-time dynamics of the system.
 In molecular dynamics, we expect to find a number of eigenvalues close to zero, followed by a spectral gap.
-These low-lying eigenvalues are indicating the number of meta stable states of the system, which are the macrostates the system stays in the longest 18 .
+These low-lying eigenvalues are indicating the number of meta-stable states of the system, which are the macro states the system stays in the longest 18 .
 We write the eigenvalue problem for the generator as
 $-\mathcal{L}\psi_{i}=\lambda_{i}\psi_{i}.$ (7)
-The eigenvalues $\lambda_{i}$ of $-\mathcal{L}$ must be non-negative, and the lowest eigenvalue $\lambda_{1}=0$ is non-degenerate under very mild conditions 36 $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq...$ We also refer to the eigenvalues as rates, and to their reciprocals as implied timescales 22
+The eigenvalues $\lambda_{i}$ of $-\mathcal{L}$ must be non-negative, and the lowest eigenvalue $\lambda_{1}=0$ is non-degenerate 33 : $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq...$ We also refer to the eigenvalues as rates, and to their reciprocals as implied timescales 17
 $t_{i}=\frac{1}{\lambda_{i}}.$ (8)
-2.4 Coarse graining
-One of the main motivations of this work is to learn an SDE representing the full dynamics (1) on a coarse-grained space.
-Coarse graining (CG) is realized by mapping the state space $\Omega$ onto a lower-dimensional space $\hat{\Omega}\subset\mathbb{R}^{d_{\xi}}$ by means of a smooth CG function $\xi$.
-We write $\nu$ for the marginal distribution of the full-space invariant measure $\mu$ projected onto CG space.
-An important object is the space $L^{2}_{\nu}$ of functions on $\hat{\Omega}$ which are square integrable with respect to the weight function $\nu$.
-The space $L^{2}_{\nu}$ is a subspace of $L^{2}_{\mu}$ comprising all functions that depend only on the value of $\xi$.
-Moreover, the orthogonal projector onto $L^{2}_{\nu}$ from $L^{2}_{\mu}$ is the conditional expectation operator 12, 13
+2.3 Coarse graining and projection
+One of the main motivations of this work is to learn an SDE representing the full dynamics (1) on a coarse grained space.
+Coarse graining (CG) is realized by mapping the state space $\Omega$ onto a lower-dimensional space $\hat{\Omega}\subset\mathbb{R}^{d}$ by means of a smooth CG function $\xi$.
+We write $\nu~{}\propto\exp(-\beta F^{\xi})$ for the marginal distribution of the full-space invariant measure $\mu$, where $F^{\xi}$ is the free energy in the CG space.
+To define dynamics in the CG space, we use the conditional expectation operator 12, 13
 $\mathcal{P}\psi(z)=\frac{1}{\nu(z)}\mathbb{E}^{\mu}[\psi(x)|\xi(x)=z],$ (9)
-where $z$ is a position in CG space. This operator calculates the average of a function $\psi$ over all $x\in\Omega$ whose projection onto CG space is the same point $z\in\hat{\Omega}$.
-Considering $\mathcal{P}$, following the exposition in 13, one can define the projected generator
+where $z$ is a position in CG space.
+This operator calculates the average of a function $\psi$ over all $x\in\Omega$ whose projection onto CG space is the same point $z\in\hat{\Omega}$.
+Following the exposition in 13 , one can define the projected generator
 $\mathcal{L}^{\xi}=\mathcal{P}\mathcal{L}\mathcal{P}.$ (10)
-It turns out its action on a function $\varphi\in L^{2}_{\nu}$ is given by
+It turns out its action on a function $\varphi=\varphi(z)$ in CG space is given by
 $\mathcal{L}^{\xi}(\phi)=\mathcal{P}[\mathcal{L}\xi]\cdot\nabla_{z}\phi+\frac{1}{2}\mathcal{P}[\nabla\xi^{T}a\nabla\xi]:\nabla_{z}^{2}\phi.$ (11)
@@
 and the practical task of coarse graining is to approximate them numerically.
-An important result, shown in 13 , and that we will use later on, is that the inner product of the projected generator with respect to the invariant measure $\nu$ in CG space is equal to the inner product of the original generator with respect to the invariant measure $\mu$ of the full space
-$\left\langle\mathcal{L}\psi\circ\xi,\,\varphi\circ\xi\right\rangle_{\mu}=\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}$ (15)
-where $\psi$ and $\varphi$ are functions in $L^{2}_{\nu}$.
-2.5 Dynamic mode decomposition and its extension
-Dynamic Mode Decomposition (DMD) is a conceptually simple method that provides a linear map to propagate the state of a dynamical system in time.
-The method was originally proposed in Ref.
-37 and since then, many variations and extensions have been proposed, see Ref.
-38, 39 for an overview.
-Basically, DMD looks for the best-fit linear operator $\mathbf{A}$ that advances the system state by solving the minimization problem in Equation (16)
-$\mathbf{A}=\operatorname*{arg\,min}_{\mathbf{A}^{\prime}}\|\mathbf{X}^{\tau}-\mathbf{A}\mathbf{X}\|_{F}=\mathbf{X}^{\tau}\mathbf{X}^{\dagger}$ (16)
-where $\mathbf{X}$ and $\mathbf{X}^{\tau}$ are collections of time-shifted snapshots of the dynamics, and $\dagger$ denotes the pseudo-inverse.
-Although DMD is a popular and widely-used method in data-driven modeling, it can only provide a linear model and the systems we are interested in for molecular dynamics are highly nonlinear.
-Improved results can be obtained by using nonlinear functions instead of just the raw state variables.
-This method is known as Extended DMD (EDMD) 40 , that can provide more accurate models by capturing nonlinearities.
-After defining a dictionary of observables, $\{\psi_{i}\}_{i=1}^{n}$ (that are functions of the state space $\Omega$), we compile the evaluations of all dictionary functions at all snapshots into matrices
-$\mathbf{\Psi}$ $=\left[\psi_{i}(x_{l})\right]_{i,l}\in\mathbb{R}^{n\times m},$ $\mathbf{\Psi}^{\tau}$ $=\left[\psi_{i}(x_{l+1})\right]_{i,l}\in\mathbb{R}^{n\times m},$
-where we assume for simplicity that subsequent snapshots are separated by time $\tau$. We then solve the optimization problem:
-$\mathbf{K}=\operatorname*{arg\,min}_{\mathbf{K^{\prime}}}\|\mathbf{\Psi}^{\tau}-\mathbf{K}^{\prime}\mathbf{\Psi}\|_{F}=\mathbf{\Psi}^{\tau}\mathbf{\Psi}^{\dagger}.$ (17)
-The solution of Equation (17) can be obtained numerically using singular value decomposition.
-In the limit of infinite data, EDMD constitutes a finite-dimensional projection (Galerkin projection) of the Koopman operator onto the subspace spanned by the dictionary functions 40 .
 2.4 Generator EDMD
-The infinitesimal generator of the Koopman operator can also be approximated by an algorithm similar to EDMD, called gEDMD 29 .
-Once again, given a set of scalar basis functions $\psi(x)=\{\psi_{1}(x),...,\psi_{n}(x)\}$, and training data $\{x_{l}\}_{l=1}^{m}$, we form the matrices
+Numerical approximations to the infinitesimal generator $\mathcal{L}$ can be obtained by a data-driven learning method called generator extended dynamic mode decomposition  28(gEDMD).
+Given a finite set of scalar basis functions $\psi(x)=\{\psi_{1}(x),...,\psi_{n}(x)\}$, and training data $\{x_{l}\}_{l=1}^{m}$ sampled from the invariant measure $\mu$, we form the matrices
 $\mathbf{\Psi}$ $=\left[\psi_{i}(x_{l})\right]_{i,l},$ $\mathcal{L}\mathbf{\Psi}$ $=\left[\mathcal{L}\psi_{i}(x_{l})\right]_{i,l},$
-using the analytical formula (7) to evaluate the second of these matrices. Solving an analogous regression problem to (17) leads to the matrix
-$\mathbf{L}^{\top}=(\mathcal{L}\mathbf{\Psi})\mathbf{\Psi}^{\dagger}=\hat{\mathbf{A}}\hat{\mathbf{G}}^{-1},$ (18)
+using the analytical formula (6) to evaluate the second of these matrices. The solution of a linear regression problem leads to the matrix approximation
+$\mathbf{L}=\hat{\mathbf{G}}^{-1}\hat{\mathbf{A}},$ (14)
 where
-$\begin{gathered}\hat{\mathbf{A}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}(\mathcal{L}\psi_{i})(x_{k})\psi_{j}(x_{k}),\\ \hat{\mathbf{G}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\psi_{j}(x_{k}).\end{gathered}$ (19)
-As for EDMD, it was proved in Ref. 29 that in the infinite-data limit, gEDMD converges to the Galerkin projection of the generator onto the space spanned by the basis functions
-$\mathbf{L}^{\top}$ $=\mathbf{A}\mathbf{G}^{-1},$ $\mathbf{A}_{ij}$ $=\left\langle\mathcal{L}\psi_{i},\,\psi_{j}\right\rangle_{\mu}$ $\mathbf{G}_{ij}$ $=\left\langle\psi_{i},\,\psi_{j}\right\rangle_{\mu}.$ (20)
-The matrices $\mathbf{G}$ and $\mathbf{A}$ are often called mass and stiffness matrix, respectively.
+$\hat{\mathbf{A}}_{ij}$ $=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\mathcal{L}\psi_{j}(x_{k}),$ $\hat{\mathbf{G}}_{ij}$ $=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\psi_{j}(x_{k}).$ (15)
+These matrices are empirical estimators of the following mass, stiffness, and generator matrices
+$\mathbf{A}_{ij}$ $=\left\langle\psi_{i},\,\mathcal{L}\psi_{j}\right\rangle_{\mu}$ $\mathbf{G}_{ij}$ $=\left\langle\psi_{i},\,\psi_{j}\right\rangle_{\mu},$ $\mathbf{L}$ $=\mathbf{G}^{-1}\mathbf{A}.$ (16)
+The empirical mass matrix $\hat{\mathbf{G}}$ is often ill-conditioned. A standard approach to circumvent this is to perform a whitening transformation based on removing small eigenvalues
+$\widehat{\mathbf{G}}$ $=\mathbf{U}\Sigma\mathbf{U}^{\top},$ $\mathbf{R}$ $=\mathbf{U}\Sigma^{-0.5}\in\mathbb{R}^{n\times r},$ $\hat{\mathbf{L}}_{r}$ $=\mathbf{R}^{\top}\hat{\mathbf{A}}\hskip 2.84544pt\mathbf{R},$ (17)
+in which $r\leq n$. Here, $\mathbf{R}$ is a transformation matrix mapping the original basis to the reduced basis
+$\mathbf{h}(x)=\mathbf{R}^{\top}\psi(x).$ (18)
+Dominant eigenvalues of the generator can be computed by diagonalizing the matrix $\hat{\mathbf{L}}$ or $\hat{\mathbf{L}}_{r}$.
 For arbitrary stochastic dynamics, the computation of $\mathbf{A}$ involves a second-order differentiation as shown in Equation (6).
 However, if the stochastic dynamics are reversible, only first-order derivatives are required to compute the matrix $\mathbf{A}$, as the generator satisfies the following integration-by-parts formula
-$\mathbf{A}_{ij}=\left<\mathcal{L}\psi_{i},\psi_{j}\right>_{\mu}=-\frac{1}{2}\int\nabla\psi_{i}\sigma\sigma^{T}\nabla\psi^{T}_{j}\,\mathrm{d}\mu.$ (21)
-If the basis functions are actually defined in a CG space $\hat{\Omega}$, that is $\psi_{i}(x)=\psi_{i}(\xi(x))$, then by the chain rule the matrix $A$ can be written as
-$\begin{split}\mathbf{A}_{ij}=-\frac{1}{2}\int\nabla_{x}\psi_{i}\sigma\sigma^{T}\nabla_{x}\psi^{T}_{j}\,\mathrm{d}\mu=-\frac{1}{2}\int(\nabla_{z}\psi_{i}\nabla_{x}\xi)\sigma\sigma^{T}(\nabla_{x}\xi^{T}\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu\\ =-\frac{1}{2}\int(\nabla_{z}\psi_{i})(\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T})(\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu.\end{split}$ (22)
+$\mathbf{A}_{ij}=\left<\psi_{i},\mathcal{L}\psi_{j}\right>_{\mu}=-\frac{1}{2}\int\nabla\psi_{i}\sigma\sigma^{\top}\nabla\psi^{\top}_{j}\,\mathrm{d}\mu.$ (19)
+Importantly, if the basis functions are actually defined in a CG space $\hat{\Omega}$, that is $\psi_{i}(x)=\psi_{i}(\xi(x))$, then by the chain rule the matrix $\mathbf{A}$ can be written as
+$\begin{split}\mathbf{A}_{ij}&=-\frac{1}{2}\int\nabla_{x}\psi_{i}\sigma\sigma^{\top}\nabla_{x}\psi^{\top}_{j}\,\mathrm{d}\mu=-\frac{1}{2}\int(\nabla_{z}\psi_{i}\nabla_{x}\xi)\sigma\sigma^{\top}(\nabla_{x}\xi^{\top}\nabla_{z}\psi^{\top}_{j})\,\mathrm{d}\mu\\ &=-\frac{1}{2}\int(\nabla_{z}\psi_{i})(\nabla_{x}\xi\sigma\sigma^{\top}\nabla_{x}\xi^{\top})(\nabla_{z}\psi^{\top}_{j})\,\mathrm{d}\mu.\end{split}$ (20)
 We refer to the matrix
-$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ (23)
+$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi\sigma\sigma^{\top}\nabla_{x}\xi^{\top}\in\mathbb{R}^{d\times d}$ (21)
 as local diffusion, and note that it is independent of the basis functions. It can therefore be computed a priori in numerical calculations.
+2.5 Random Fourier features
+The gEDMD algorithm requires choosing a set of basis functions $\psi(x)$. In this work, we use random Fourier features (RFFs), which are defined as
+$\psi(x)=\{\cos(\omega_{1}^{\top}x),\,\sin(\omega_{1}^{\top}x),...,\cos(\omega_{n}^{\top}x),\,\sin(\omega_{n}^{\top}x)\}.$ (22)
+The vectors $\omega_{1},\ldots,\omega_{n}$ are random frequency vectors drawn from a spectral distribution $\rho$.
+RFFs provide a low-rank approximation to a reproducing kernel function 29 , and can therefore generate a powerful basis without the need for manual basis set design.
+The precise relation between kernel-based gEDMD and random features was presented in Ref.
+34.
+In the following applications, we use the spectral measure associated to a Gaussian squared exponential kernel with bandwidth parameter $\gamma$
+$k(x_{i},x_{j})=\exp\left(-\frac{\|x_{i}-x_{j}\|)^{2}}{2\gamma^{2}}\right),$ (23)
+or to a periodic Gaussian kernel 35 on periodic domains, such as dihedral coordinates.
 3 Methods
-3.1 Overview and Notation
-Successful application of the gEDMD method requires selecting proper and expressive basis functions.
-This is not a trivial task and can highly depend on the dynamics of interest.
-To alleviate this, use of kernel methods has been proposed in Ref.
-32 .
-The kgEDMD method is based on choosing the basis functions as features of a reproducing kernel $k(x,y)$ at the training data:
-$\Phi=\left[k(\cdot,x_{1}),k(\cdot,x_{2}),...,k(\cdot,x_{m})\right]^{\top}.$ (24)
-We consider $\Phi$ as a vector-valued function, that is, we write
-$\Phi(x)=\left[k(x,x_{1}),k(x,x_{2}),...,k(x,x_{m})\right]^{\top}$ (25)
-for the $m$-dimensional vector obtained by evaluating all the kernel features at $x$.
-Since kernels are two-argument functions, we can take their derivatives with respect to both arguments, and in fact we will need to distinguish these derivatives when formulating our methods.
-In general, we simply write $\nabla_{1}k$ and $\nabla_{2}k$ for gradients of the kernel function with respect to its first and second argument.
-Similarly, we write $\nabla_{1}\Phi,\,\nabla_{2}\Phi$ for the matrix-valued functions we obtain by assembling the gradients of all kernel features:
-$\nabla_{1}\Phi(x)$ $=\begin{bmatrix}\nabla_{1}k(x,x_{1})\\ \vdots\\ \nabla_{1}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d},$ $\nabla_{2}\Phi(x)$ $=\begin{bmatrix}\nabla_{2}k(x,x_{1})\\ \vdots\\ \nabla_{2}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d}.$ (26)
-One of the widely used kernel functions is the Gaussian kernel given by
-$k(x_{i},x_{j})=\exp\left(-\frac{\|x_{i}-x_{j}\|)^{2}}{2\sigma^{2}}\right),$ (27)
-where $\sigma$ is the standard deviation or bandwidth of the Gaussian kernel. If the state space is a periodic domain, we make use of a periodic Gaussian kernel instead 41
-$k(x_{i},x_{j})=\exp\left(-\frac{\sum_{i=1}^{d}\sin\left(\frac{\pi}{p}(x_{i}-x_{j})\right)^{2}}{2\sigma^{2}}\right),$ (28)
-where $\sigma$ is again the bandwidth of the Gaussian kernel, $p$ is the half-period of the domain, and $d$ is the dimension of system.
-3.2 Kernel-based Koopman generator EDMD
-Using the kernel feature basis as in Equation (24), we can seek a data-driven approximation to the eigenfunctions of the Koopman generator in the form $\psi_{i}(x)=\mathbf{u}_{i}^{\top}\Phi(x)$.
-As detailed in Ref.
-32 , the coefficient vectors $\mathbf{u}_{i}$, are then found as solutions of the following generalized matrix eigenvalue problem
-$\hat{\mathbf{A}}\mathbf{u}_{i}:=\frac{1}{2}\sum_{l=1}^{d}(\widehat{\mathbf{A}}^{(l)})^{\top}\widehat{\mathbf{A}}^{(l)}\mathbf{u}_{i}=\widehat{\lambda}_{i}\widehat{\mathbf{G}}\widehat{\mathbf{G}}\mathbf{u}_{i},$ (29)
-where $d$ is the dimension of the state space.
-The matrices $\hat{\mathbf{A}}^{(l)}$ and $\hat{\mathbf{G}}$ are based on the reversible formulation for gEDMD, and essentially contain pairwise evaluations of the kernel and its first-order derivatives
-$\begin{gathered}\left[\widehat{\mathbf{G}}\right]_{ns}=k(x_{n},x_{s})\\ \left[\widehat{\mathbf{A}}^{(l)}\right]_{ns}=\sigma_{l}(x_{n})^{\top}\nabla_{2}k(x_{n},x_{s}).\end{gathered}$ (30)
-The kernelized mass matrix $\hat{\mathbf{G}}$ is often highly ill-conditioned.
-Therefore, we compute a reduced generator by performing a whitening transformation based on removing small eigenvalues of the mass matrix
-$\widehat{\mathbf{G}}$ $=\mathbf{U}\Sigma\mathbf{U}^{T},$ $\mathbf{R}$ $=\mathbf{U}\Sigma^{-0.5}\in\mathbb{R}^{m\times r},$ $\hat{\mathbf{L}}_{r}$ $=\mathbf{R}^{T}\hat{\mathbf{A}}\hskip 2.84544pt\mathbf{R},$ (31)
-in which $r\leq m$. Here, $\mathbf{R}$ is a transformation matrix mapping the original basis built by the feature map $\Phi$, to the reduced basis
-$\mathbf{h}(x)=\mathbf{R}^{\top}\Phi(x)\in\mathbb{R}^{r}.$ (32)
-In preparation for later, we also introduce the notation
-$\nabla\mathbf{h}(x)=\left[\partial_{j}h_{l}(x)\right]_{l,j}=\mathbf{R}^{\top}\nabla_{1}\Phi(x)\in\mathbb{R}^{r\times d}$ (33)
-for the gradients of all reduced features evaluated at point $x$.
-Diagonalization of the reduced generator matrix $\hat{\mathbf{L}}_{r}$ then provides the final estimates for the generator’s eigenvalues and eigenfunctions.
-3.3 System identification
-The reduced generator matrix $\hat{\mathbf{L}}_{r}$ provides an empirical model for the Koopman generator $\mathcal{L}$ on the space of whitened and rank-reduced features $\mathbf{h}$.
-Hence we can use this model towards other ends besides identification of the meta stable sets.
-Among other purposes, we can use it to identify an effective SDE governing the CG dynamics as in Equation (13), using a procedure that was already sketched in Ref.
-29 .
-We parameterize the CG dynamics as a reversible SDE, which is justified if the original dynamics are reversible 13 .
-As discussed in Section 2 and in Ref.
-33 , reversible systems can be characterized by the diffusion $a^{\xi}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ and a scalar potential $F^{\xi}:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}$, from which the drift is then obtained by:
+The main proposal in this work is a method to identify a diffusion field for the coarse-grained dynamics based on data-driven approximations to the generator.
+We also show how the quality of the diffusion field in terms of reproducing dominant eigenvalues can be assessed, and finally, how the diffusion can be complemented to identify the complete CG dynamics.
+We recall that the dynamical equation in CG space is given by (12), where the drift can be written as follows because of reversibility 30
 $b^{\xi}=-\dfrac{1}{2}a^{\xi}\nabla_{z}F^{\xi}+\dfrac{1}{2}\nabla\cdot a^{\xi}.$ (24)
-Thus, to identify the SDE, we need to learn these two components, which can be broken up into two separate problems.
-From now on, we employ a kernel $k$ on the CG space $\hat{\Omega}$, that is $k=k(z,z^{\prime})$.
-3.3.1 Force matching
-Knowing the full-process scalar potential field $F$ and the CG function $\xi$, the effective force can be obtained by a well-known technique called force matching 15 via the following minimization problem
-$\nabla_{z}F^{\xi}=\underset{g\in L^{2}_{\nu}}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|g(\xi(x_{i}))-f^{\xi}_{\text{lmf}}(x_{i})\right\|^{2},$ (35)
-where $m$ is the data size and $f^{\xi}_{\text{lmf}}$ is called local mean force and defined as follows
-$\begin{gathered}f^{\xi}_{\text{lmf}}=-\nabla_{x}F\cdot G^{\xi}+\nabla_{x}\cdot G^{\xi}\\ G^{\xi}=\nabla_{x}\xi[(\nabla_{x}\xi)^{T}\nabla_{x}\xi]^{-1}.\\ \end{gathered}$ (36)
-In line with the kernel-based approach we follow in this work, we parameterize the optimal vector field $g$ as the gradient of a linear combination of the feature basis $\Phi$ in Equation (24)
-$F^{\xi}(z)$ $=\alpha^{\top}\Phi(z),$ $g(z)$ $=\alpha^{\top}\nabla_{1}\Phi(z)=\sum_{i=1}^{m}\alpha_{i}\nabla_{1}k(z,z_{i}).$ (37)
-The force matching problem thus breaks down to determination of $m$ optimal expansions coefficients $\alpha$.
-We stress, however, that any other parameterization could be used for this step, for example a deep learning-based method 11 .
-3.3.2 Kernel-based diffusion learning
-Our main contribution is a learning method for the effective diffusion, complementing the effective potential $F^{\xi}$.
-We first use training data of the full space dynamics and the kernel $k$ on the CG space $\hat{\Omega}$, to learn a reduced matrix representation $\hat{\mathbf{L}}_{r}$ of the full generator.
-Recalling Equation (15), the matrix $\hat{\mathbf{L}}_{r}$ is at the same time an approximation to the CG generator $\mathcal{L}^{\xi}$.
-On the CG space, the reversible generator $\mathcal{L}^{\xi}$ also satisfies the integration-by-parts formula (21),
-$\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}=-\frac{1}{2}\int_{\hat{\Omega}}\nabla_{z}\psi(z)a^{\xi}(z)\nabla_{z}\varphi(z)^{\top}\,\mathrm{d}\nu(z),$ (38)
-which only features an explicit dependence on the unknown effective diffusion $a^{\xi}$, see Equation (14).
-Using the reduced basis $\mathbf{h}(z)=\mathbf{R}^{\top}\Phi(z)$, we parameterize the effective diffusion as
-$\sigma^{\xi}_{\alpha}(z)=\sum^{r}_{l=1}\alpha_{l}h_{l}(z),$ (39)
-where, in the most general case, $\alpha_{l}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ must be matrix-valued coefficients.
-To arrive at a compact expression, we gather all unknown expansion coefficients $\alpha_{l}$ as a third order tensor $\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}$, which allows us to write
-$\sigma^{\xi}_{\alpha}(z)=\alpha\cdot_{|3,1}\mathbf{h}(z)=\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (40)
-using the notation $\mathbf{a}\cdot_{|i,j}\mathbf{b}$ to denote contraction of two tensors $\mathbf{a},\,\mathbf{b}$ over indices $i$ and $j$ of commensurate dimension.
-Plugging this into Equation (38) and using the data-driven estimator of Equation (30), we obtain a parametric model for the reduced generator matrix $\hat{\mathbf{L}}_{r}$
-$\begin{split}\left[\hat{\mathbf{L}}^{\xi}_{\alpha}\right]&=\dfrac{1}{2m}\sum_{k=1}^{m}\nabla\mathbf{h}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})^{\top}\nabla\mathbf{h}^{\top}(z_{k})\\ &=\dfrac{1}{2m}\sum_{k=1}^{m}\mathbf{S}^{\alpha}(z_{k})\mathbf{S}^{\alpha}(z_{k})^{\top},\\ \mathbf{S}^{\alpha}(z)&:=\nabla\mathbf{h}(z)\sigma_{\alpha}^{\xi}(z).\end{split}$ (41)
-Combining Equations (33) and (40), we can compactly write
-$\mathbf{S}^{\alpha}(z)=\mathbf{R}^{\top}\nabla_{1}\Phi(z)\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (42)
-which is now expressed solely in terms of the kernel features $\Phi$, the whitening transformation $\mathbf{R}$, and the unknown coefficient tensor $\alpha$.
-We then learn the effective diffusion by matching the parameterized generator matrix $\hat{\mathbf{L}}^{\xi}_{\alpha}$ to the true reduced generator matrix $\hat{\mathbf{L}}_{r}$, therefore solving the following minimization problem:
-$\alpha^{*}=\underset{\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}}{\arg\min}\left[\left\|\hat{\mathbf{L}}_{r}-\hat{\mathbf{L}}^{\xi}_{\alpha}\right\|_{2}^{2}+\gamma\left\|\frac{\partial\sigma}{\partial z}\right\|_{2}^{2}\right],$ (43)
-where $\hat{\mathbf{L}}^{\xi}_{\alpha}$ is calculated from $\alpha$ via Equations (41) and (42).
-To avoid over-fitting and to suppress oscillations in $\sigma$, we found that introducing the second term to regularize the derivatives of the diffusion provides the best results.
-Here, the effective diffusion is represented by a full matrix.
-However, for simplicity, we only learn a diagonal diffusion matrix for the numerical results reported in the following section.
-The method can be summarized in Algorithm 1.
-Algorithm 1 Learning Effective Dynamics with kgEDMD
-Input: full space data $\{x_{k}\}_{k=1}^{m}$ in $\mathbb{R}^{d}$, CG map $\xi$, kernel function $k$,
-Truncation rule for $r$ in Equation (31), regularization $\gamma$
+3.1 Diffusion Learning
+By Equation (13), the analytical effective diffusion $a^{\xi}$ is the best-approximation of the local diffusion $a^{\xi}_{\mathrm{loc}}$ by a (matrix-valued) function on the CG space.
+Hence, we can solve the following data-based minimization problem
+$a^{\xi}=\underset{a=a(z)}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|a(\xi(x_{i}))-a^{\xi}_{\mathrm{loc}}(x_{i})\right\|_{F}^{2},$ (25)
+where $\|\cdot\|_{F}$ is the Frobenius norm for matrices. We parametrize the diffusion field $a^{\xi}$ element-wise as a linear combination of the reduced RFF basis
+$(a^{\xi}_{\alpha})_{ij}(z)=\sum_{l=1}^{r}\alpha^{ij}_{l}h_{l}(z)=\alpha\cdot_{|3,1}\mathbf{h}(z)=\alpha\cdot_{|3,1}\mathbf{R}^{\top}\psi(z),$ (26)
+where we view the coefficient array $\alpha$ as a third-order tensor of dimension $d\times d\times r$, and the symbol $\cdot_{|i,j}$ denotes contraction over indices $i$ and $j$ of two arrays.
+The parametrization must be symmetric, i.e.
+$a^{\xi}_{ij}=a^{\xi}_{ji}$, and we may also choose to set specific elements to zero, for example to enforce a diagonal diffusion field.
+With the parametrization (26), the minimization problem (25) becomes a linear regression problem that can be directly solved, potentially after regularization.
+3.2 Recovery of Spectral Properties
+After solving the minimzation problem (25), we can make use of the gEDMD method to assess the dynamical properties of the learned SDE in CG space.
+Using the integration-by-parts formula (19), the elements of the reduced generator matrix corresponding to the diffusion field (26) with coefficient array $\alpha$ are
+$\left\langle h_{r},\,\mathcal{L}^{\alpha}h_{s}\right\rangle_{\nu}=-\frac{1}{2}\int\nabla h_{r}(z)a^{\xi}_{\alpha}(z)\nabla h_{s}(z)^{\top}.$
+In matrix notation, this leads to the following explicit formula for the parametrized generator matrix, which can be computed directly without resorting to numerical simulations of the CG dynamics
+$\hat{\mathbf{L}}_{r}^{\alpha}=\sum_{i=1}^{m}\mathbf{R}^{\top}\nabla_{z}\psi(z_{i})\left[\alpha\cdot_{|3,1}\mathbf{R}^{\top}\psi(z_{i})\right]\nabla_{z}\psi(z_{i})^{\top}\mathbf{R}.$ (27)
+Properties inferred from the matrix $\hat{\mathbf{L}}^{\alpha}_{r}$ can be compared to those obtained from the original gEDMD matrix $\hat{\mathbf{L}}_{r}$ estimated off the full-space simulation data.
+For example, diagonalization of both $\hat{\mathbf{L}}_{r}$ and $\hat{\mathbf{L}}^{\alpha}_{r}$ leads to estimates $\lambda_{i}$ and $\lambda_{i}^{\alpha}$ for the dominant generator eigenvalues, which can be systematically compared.
+We mainly resort to comparing dominant eigenvalues in the examples below, but we point out that a more detailed assessment is possible: for instance, by computing matrix exponentials $\exp\left(t\hat{\mathbf{L}}_{r}\right)$ and $\exp\left(t\hat{\mathbf{L}}^{\alpha}_{r}\right)$, time-correlation functions can also be evaluated.
+3.3 Learning the Effective Potential
+We have seen that the accuracy of the effective diffusion field largely determines the dynamical properties of the coarse grained dynamics.
+In order to run simulations of the CG dynamics, and to ensure thermodynamic consistency, the effective potential $F^{\xi}$ also must be learned in a parametric form.
+This is not the main focus of our study, hence we just point out a few options.
+A well-known and generally applicable technique is force matching 15 , which is based on the following minimization problem for the effective force
+$\nabla_{z}F^{\xi}=\underset{g=g(z)}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|g(\xi(x_{i}))-f^{\xi}_{\text{lmf}}(x_{i})\right\|^{2},$ (28)
+where $f^{\xi}_{\text{lmf}}$ is called local mean force and defined as follows
+$f^{\xi}_{\text{lmf}}$ $=-\nabla_{x}F\cdot G^{\xi}+\nabla_{x}\cdot G^{\xi},$ $G^{\xi}$ $=\nabla_{x}\xi[(\nabla_{x}\xi)^{\top}\nabla_{x}\xi]^{-1}.$ (29)
+We point out the similarity to (25), which also led us to the name local diffusion for $a^{\xi}_{\mathrm{loc}}$.
+The effective potential can be parametrized as a linear combination of basis functions, such as random features, or as a deep neural network 11 .
+In low-dimensional CG spaces, it also possible to approximate the projected invariant distribution $\nu$ as a linear combination of kernel functions centered at the data sites, known as kernel density estimate (KDE) 36 .
+Since we only consider low-dimensional CG spaces here, we opt for the KDE option in the examples below.
+Algorithm 1 Learning Effective Dynamics
+Input: full space data $\{x_{k}\}_{k=1}^{m}$ in $\mathbb{R}^{d}$, CG map $\xi$, kernel function with spectral measure $\rho$,
+Truncation rule for $r$ in Equation (17), regularization parameters
 1:Diffusion learning
 2:     Compute local diffusion $a^{\xi}_{\rm loc}$ as in Equation (21).
-3:     Form mass and stiffness matrices $\mathbf{G}$ and $\mathbf{A}^{(l)}$ as in Equation (30)
-4:     Compute reduced generator $\hat{\mathbf{L}}_{r}$ as in Equation (31)
-5:     Perform the minimization problem as in Equation (43)
-6:     Form the effective diffusion $\sigma^{\xi}_{\alpha}$ as in Equation (39)
-7:Force matching
-8:     Compute the local mean force $f^{\xi}_{\text{lmf}}$ as in Equation (36)
-9:     Perform the minimization problem as in Equation (35)
-10:     Form the effective potential $F^{\xi}$ as in Equation (37)
+3:     Generate random feature basis $\psi$ as in Equation (22).
+4:     Compute reduced basis $\mathbf{h}$ according to (18).
+5:     Compute reduced generator matrix $\hat{\mathbf{L}}_{r}$ as in Equation (17).
+6:     Perform the minimization problem as in Equation (25).
+7:     Compute learned generator matrix $\hat{\mathbf{L}}_{r}^{\alpha}$ by (27) and compare its properties to $\hat{\mathbf{L}}_{r}$.
+8:Learning the full CG dynamics
+9:     Learn effective potential $F^{\xi}$ by KDE or force matching (28).
+10:     Effective drift is given by (24).
+3.4 Overdamped Models for Molecular Systems
+For the molecular examples considered in Sections 4.2 and 4.3, the dynamics are not governed by the reversible overdamped dynamics (3).
+Nevertheless, it can be expected that their position space dynamics are similar to the overdampled process after a re-scaling of time.
+This can be rigorously shown for the underdamped dynamics used in Section 4.2 37 .
+Fitting an overdamped model is simpler than modeling the full non-reversible dynamics if a non-linear CG map is used.
+Therefore, we compute the local diffusion as
+$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi(x)\frac{2}{\beta\gamma}M^{-1}\nabla_{x}\xi^{\top}(x),$ (30)
+where $M$ is the diagonal mass matrix of all atoms, $\beta$ is the inverse temperature, $\gamma$ is the friction, and $\nabla_{x}\xi(x)$ is the Jacobian of the CG map.
+This way of computing the local diffusion corresponds to a reversible overdamped Langevin process in full position space, meaning that the resulting CG model is an approximation to that same overdamped process.
+This also means that the eigenvalues and timescales of the overdamped model may differ from the ones obtained by directly approximating the full non-reversible process, but the meta-stable states will remain the same.
 4 Examples
-To show the effectiveness of the proposed method, we apply it to a two-dimensional model system defined by the Lemon-slice potential, and to MD simulation data of the alanine dipeptide, which is a widely-used test case in molecular dynamics.
-As the optimization block for diffusion learning is non-convex, we use a library function from scipy’s optimization toolbox, but solve the problem repeatedly for $50$ randomly generated initial guesses.
-The optimization block for the force matching, on the other hand, is done via ridge regression.
+To show the effectiveness of the proposed method, we apply it to a two-dimensional model system defined by the Lemon-slice potential, and to MD simulation data of the alanine dipeptide and of the mini-protein Chignolin, which are widely-used test cases in molecular dynamics.
+The optimization problems for learning the effective force and diffusion are solved by standard minimizers from scipy’s optimization toolbox.
+Note that in case the sought after object in the optimization blocks are decoupled, e.g.
+under the assumption that the effective diffusion is diagonal, we replace the minimizers with a simple least-square problem, making the step much cheaper.
 4.1 Lemon-Slice potential
@@
 We further down sample them to $1000$ samples each for learning effective force and diffusion.
-As shown in previous studies, the polar angle $\phi$ is a suitable CG coordinate for this system, as it resolves all four meta stable states
+As shown in previous studies, the polar angle $\phi$ is a suitable CG coordinate for this system, as it resolves all four meta-stable states
 $\xi(x,y)=\phi.$ (33)
-For this system, analytical expressions for the effective drift and diffusion along $\xi$ can be obtained by a slight modification of the results in 42 , and serve as reference values.
-We use the kgEDMD algorithm on the reaction coordinate $\xi$, to identify the generator eigenvalues and meta stable states and, subsequently, to identify an effective dynamics along $\xi$ using Algorithm 1.
-As a first try, we consider the Gaussian kernel for Galerkin approximation of the generator.
-Notice that the CG dynamics of Lemon-slice problem is periodic in the polar angle, however, a generic Gaussian kernel is not so.
-Thus, it cannot capture the periodicity of the system of interest.
-As an alternative, we also use the periodic Gaussian kernel as in Equation (28) and analyze both.
-The kernel bandwidth in either versions of Gaussian kernel is optimized using cross validation based on the VAMP-score 27 .
+For this system, analytical expressions for the effective drift and diffusion along $\xi$ can be obtained by a slight modification of the results in Ref. 38 , and serve as reference values.
+We apply our learning method with random Fourier features on the reaction coordinate $\xi$, to identify the generator eigenvalues and meta-stable states and, subsequently, to identify an effective dynamics along $\xi$ using Algorithm 1.
+As the polar angle is a periodic reaction coordinate (RC), we use the spectral measures associated to both a periodic and non-period Gaussian kernel and compare them.
+The kernel bandwidth in either versions of Gaussian kernel is optimized using cross validation based on the VAMP-score 26 .
 Details on the VAMP-score analysis are reported in the appendix.
 4.1.2 Meta-stability analysis
-Figure 1 c shows the leading eigenvalues obtained from the kgEDMD model.
+Figure 1c shows the leading eigenvalues obtained from the generator matrix $\hat{\mathbf{L}}_{r}$.
 As one notices, there are four dominant eigenvalues followed by a gap.
 These four eigenvalues are corresponding to the four minima in the potential field.
-Having determined the eigenvectors of the generator, we can perform robust Perron Cluster Cluster Analysis (PCCA+) 43 algorithm to assign to each sample point its membership to each meta stable state.
+Having determined the eigenvectors of the generator, we can perform robust Perron Cluster Cluster Analysis (PCCA+) 39 algorithm to assign to each sample point its membership to each meta-stable state.
 Figure 1b shows that the four potential minima are perfectly recovered in this way.
@@
 The relative error of these eigenvalues compared to the reference is shown in (d).
-A comparison of the leading eigenvalues of the reference model $\hat{\mathbf{L}}_{r}$ and the optimal matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$ is shown in Figure 1 c.
+A comparison of the leading eigenvalues of the reference model $\hat{\mathbf{L}}_{r}$ and the learned matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$ for the optimal parameters $\alpha$ is shown in Figure 1c.
 Both choices of the kernel function lead to satisfactory results, the periodic kernel provides slightly higher accuracy in approximation of the generator eigenvalues.
 Note that the kernel bandwidth is tuned for each kernel function separately.
-4.1.3 Identification of CG dynamics
+4.1.3 Analysis of the CG dynamics
 The learned generator providing the eigenvalues reported above is built upon the effective diffusion shown in Figure 2b, which is almost perfectly following the reference.
@@
 From the effective force and diffusion, the effective drift can be obtained according to Equation (24), which is also compared against the analytical expression in Figure 2c, likewise showing very good agreement.
-Having effective drift and diffusion, we are able to build the learned SDE governing the CG coordinate.
+With the effective drift and diffusion fields, we are able to simulate the learned SDE governing the CG coordinate.
 We use the Euler-Maruyama scheme to integrate the learned and reference SDEs with integration time step of $dt=10^{-3}$.
 Figure 2d shows two trajectories of the CG coordinate $\phi$ for both dynamics for $10^{4}$ time steps, using the same Brownian motion for both trajectories.
-The propagated learned system follows the reference closely, with both systems staying long times in each meta stable state, and rarely swapping in between those.
-The results demonstrate that the proposed method can approximate the full system’s meta stable sets well, and identify a suitable SDE for CG dynamics which is accurate even on the level of individual trajectories.
+The propagated learned system follows the reference closely, with both systems staying long times in each meta-stable state, and rarely swapping in between those.
+Combined, the results above demonstrate that the proposed method can approximate the full system’s meta-stable sets well, and identify a suitable SDE for CG dynamics which is accurate even on the level of individual trajectories.
 Figure 2: Application of Algorithm 1 to identify angular dynamics for the Lemon-slice system.
 Effective force in (a), effective diffusion in (b), effective drift in (c), and integration of an example trajectory, using both the reference and learned SDE in (d).
-In this example, both the full dynamics and the learned effective dynamics possess state-dependent diffusion fields.
-We compare the properties of the learned CG model with variable diffusion to those of a CG dynamics with constant diffusion.
-For the latter, we set the effective diffusion constant to $a=\frac{2}{\beta}=2$.
-We propagate the corresponding SDEs for a sufficiently large span of time, and estimate a new generator EDMD model based on these simulations, this time using Random Fourier Features as observables 44 .
+As a final analysis, we compare the properties of the learned CG model with variable diffusion to those of a CG dynamics with constant diffusion, in order to demonstrate the necessity of allowing a state-dependent diffusion.
+We set the effective diffusion for the constant model to $a=\frac{2}{\beta}=2$.
+We propagate the corresponding SDEs for a sufficiently large span of time, and estimate a new generator EDMD model based on these simulations.
 Figure 3, shows the eigenvalues of the generator for these cases compared to the learned generator built upon the original dataset.
 The result shows that learning a state-dependent diffusion is necessary to recover the original system’s leading eigenvalues.
-To show the scalability of the method, we repeat the analysis on a real system in the following section.
 Figure 3: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (SDD, orange) and with constant diffusion (CD, green).
-As a comparison, we show the eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ using the original dataset (blue triangles).
+As a comparison, we show the eigenvalues of the generator $\hat{\mathbf{L}}_{r}$ using the original dataset (blue).
 Note that the first eigenvalue is omitted as it is zero.
@@
 4.2.1 System introduction
+Alanine dipeptide is a model system widely used in method development for simulation studies of macro-molecules.
+Figure 4 shows the graphical representation of Alanine dipeptide.
+It is well known that the dynamical behavior of the molecule can be expressed in terms of the backbone dihedral angles $\phi$ and $\psi$, which constitute the two-dimensional reaction coordinate space defining the CG map $\xi$:
+$\xi(x)=\begin{bmatrix}\phi(x)&\psi(x)\end{bmatrix}.$ (34)
 […]
-$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi(x)\frac{2}{\beta\gamma}M^{-1}\nabla_{x}\xi^{\top}(x)$ (47)
-where $M$ is the diagonal mass matrix of all atoms, $\beta$ is the inverse temperature, $\gamma$ is the friction, and $\nabla_{x}\xi(x)$ is the Jacobian of dihedral angles.
-This way of computing the local diffusion corresponds to a reversible overdamped Langevin process in full position space, meaning that the resulting CG model is an approximation to that same overdamped process.
-Fitting an overdamped model is simpler than modeling the full non-reversible underdamped Langevin process if a non-linear CG map is used.
-In addition, the underdamped process is expected to be close to an overdamped model in position space after a re-scaling of time.
-However, this also means that the eigenvalues and timescales of the overdamped model are expected to differ from the ones obtained by directly approximating the underdamped Langevin model, while the meta stable states will remain the same.
-The free energy landscape of the system with respect to these two angles is shown in Figure 4.
+The familiar free energy landscape of the system with respect to these two angles is shown in Figure 4, displaying four minima, two on the left side, and two in the central part.
 Figure 4: Graphical representation of the alanine dipeptide molecule on the left, and the reference free energy profile in two-dimensional dihedral angle space on the right.
-The free energy landscape shows four minima, two on the left side, and two in the central part.
-We apply the kgEDMD algorithm to find the meta stable sets, and then use Algorithm 1 to learn the effective force and a state-dependent effective diffusion field in the dihedral angle space.
-Because of the periodicity of the CG coordinates, $\phi$ and $\psi$, we adopt the periodic Gaussian kernel.
-Similar to the previous example, we tune the bandwidth in the kernel function using the VAMP-score.
-4.2.2 Meta stability analysis
-Figure 5 a shows the leading eigenvalues of the generator obtained from the kgEDMD model.
+We apply the gEDMD algorithm with random Fourier features to find the meta-stable sets, and then use Algorithm 1 to learn the effective force and a state-dependent effective diffusion field in the dihedral angle space.
+Because of the periodicity of the CG coordinates, $\phi$ and $\psi$, the spectral measure corresponds to a periodic Gaussian kernel.
+Similar to the previous example, we tune the bandwidth of the kernel function using the VAMP-score.
+4.2.2 Meta-stability analysis
+Figure 5a shows the leading finite timescales by taking reciprocals of the first three nonzero eigenvalues of the generator obtained from the gEDMD matrix $\hat{\mathbf{L}}_{r}$ (error bars in the figure are generated by analyzing $5$ independent subsampled sets of the original data set, each comprising $50000$ samples).
 Figure 5: Approximation of generator for alanine dipeptide.
-The dominant eigenvalues of the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error of these eigenvalues compared to the reference is shown on the right.
-The figure indicates the four dominant eigenvalues which are corresponding to the four minima in the free energy landscape followed by a gap.
-In addition, we also show the eigenvalues of the generator $\mathbf{L}^{\xi}_{\alpha}$ corresponding to the optimal effective diffusion, which agree well with the reference.
-Errorbars in the figures are generated by analyzing $5$ independent subsampled sets of the original data set, each one comprising $2000$ samples.
-4.2.3 Identification of CG dynamics
-For this 2-dimensional CG dynamics, we can express the diffusion field in a $2\times 2$ full matrix.
-Here we assume though, that the learned diffusion is a diagonal matrix.
-Figure 6 shows the first and second diagonal terms of the learned diffusion field.
-Figure 6: The first (a) and second (b) diagonal terms of the learned diffusion covariance matrix $a$, the reference free energy surface (c) and the free energy surface learned via kernel-based force matching (d).
-For the diffusion learning process, we incorporate $2000$ samples of the available dataset.
-The next step is to apply force matching algorithm to obtain the effective force.
-The reference and learned effective free energy surface are depicted in Figures 6c and  6d, respectively.
-It it noticeable that the learned free energy surface correctly captures all energetic minima, and approximates the energy barriers properly.
-We also note that $m=50,000$ samples were required to obtain the results in Figure 6 d with kernel-based force matching.
+The dominant timescales corresponding to the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error of these timescales compared to the reference is shown on the right.
+The figure indicates the three dominant timescales which are corresponding to the four minima in the free energy landscape followed by a gap.
+In addition, we also show the timescales corresponding to the generator $\mathbf{L}^{\xi}_{\alpha}$ based on the optimal effective diffusion, which agree well with the reference.
+As another way for validation of the timescales, we also computed the leading implied timescales using an MSM trained on the original simulation data, which does not rely on the overdamped assumption.
+These timescales exceed those of the generator model $\mathbf{L}^{\xi}_{\alpha}$ by a uniform factor of about $100$, their re-scaled values are shown as black dashed lines.
+Thus, we observe that the dynamics in CG space based on the overdamped assumption is accelerated by a factor $100$ for this example, which confirms similar observations from previous studies.
+In principle, the exact transition timescales of the full simulation can be recovered by re-scaling the friction term of the CG model accordingly.
+4.2.3 Analysis of the CG dynamics
+For this 2-dimensional coarse graining, we can express the diffusion field as a $2\times 2$ full matrix.
+For simplicity, however, we assume that the learned diffusion is a diagonal matrix.
+Figure 6 shows the first and second diagonal terms of the learned diffusion field based on $50000$ samples of the available dataset.
+Figure 6: The first (a) and second (b) diagonal terms of the learned diffusion covariance matrix $a$, the reference free energy surface (c) and the free energy surface learned via KDE (d).
+To learn the effective potential, we found that the KDE method works best.
+The reference and learned effective free energy surfaces are depicted in Figures 6c and  6d, respectively.
+It it noticeable that the learned free energy surface correctly captures all energetic minima and barriers up to some minor spurious behaviour close to the transition regions.
+We emphasize once again that this approximation could probably be improved further by using a more accurate learning method.
 […]
+Figure 7: Left: Free energy surface learned via KDE. Right: estimated free surface from histogramming the simulated CG dynamics.
 […]
+Figure 8: Kinetic consistency of the CG dynamics for alanine dipeptide.
+Left: PCCA+ membership analysis applied to simulation data of the CG dynamics.
+Right: slowest finite timescales calculated using an approximation of the generator from the reference dataset (blue) and the propagated CG dynamics with state-dependent diffusion (SDD, orange) as well as constant diffusion (CD, green), compared to those obtained via a Markov state model (black).
+4.3 Chignolin
+4.3.1 System introduction
 […]
-Figure 8: Dominant eigenvalues of the generator, using models built on simulation data of the learned coarse-grained dynamics with state-dependent diffusion (Prop.-SDD, green) and with constant diffusion (Prop.-CD, red).
-As a comparison, we show the eigenvalues of the optimal learned generator matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$, assuming either state-dependent diffusion (blue dots) or constant diffusion (orange triangles).
-Note that the first eigenvalue is omitted as it is zero.
+$\xi(x)=\begin{bmatrix}\mathrm{TIC}_{1}(x)&\mathrm{TIC}_{2}(x)\end{bmatrix}.$ (35)
+By projecting the atomistic positional information of the system onto this $2$-dimensional TICA space and computing the histogram of the data, the free energy surface can be obtained, as shown in Figure 2.
+Figure 9: Graphical representation of CLN025 on the left  [^2] The image is generated using Protein Data Bank in Europe platform., and the reference free energy surface in the two-dimensional TICA space on the right.
+The left-hand side minimum corresponds to the folded state, the bottom right minimum corresponds to the unfolded state and the top one associates to the misfolded state.
+The free energy surface shows three minima, representing the three conformational states of folded, unfolded and misfolded. We perform the same analysis as for the previous example.
+4.3.2 Meta-stability analysis
+To find the timescales of the system, we applied the gEDMD method with random Fourier features as before, and computed the eigenvalues of the generator model $\hat{\mathbf{L}}_{r}$.
+Figure 10 shows the corresponding timescales of the system, which are the inverse of the generator’s eigenvalues.
+The figure indicates the two leading timescales of the system (the first one associated to the stationary distribution is not shown) corresponding to the three meta-stable sets, followed by a spectral gap.
+Moreover, we show that the timescales of the CG generator $\mathbf{L}^{\xi}_{\alpha}$ for the optimal effective diffusion are very similar, the relative errors shown on the right of the same figure are sufficiently small.
+Also, we observe that the gEDMD timescales are once again uniformly re-scaled compared to the leading timescales of an MSM estimated on the original data, which is not subject to the overdamped assumption.
+The re-scaling factor is quite drastic this time, reducing micro-second timescales of the full system to less than pico-seconds for the CG dynamics.
+Nevertheless, as the re-scaling is again uniform, the original timescales can in principle be recovered by re-scaling the friction term.
+Error-bar figures were again generated by analyzing $5$ independent subsampled sets, each comprising $1.6\times 10^{5}$ samples.
+Figure 10: Approximation of generator for Chignolin.
+The slowest finite timescales corresponding to the reference generator $\hat{\mathbf{L}}_{r}$ and the learned generator $\hat{\mathbf{L}}_{\alpha}^{\xi}$ built upon the learned effective diffusion on the left, and the relative error on the right.
+The black dashed lines on the left indicate re-scaled timescales obtained from an MSM on the original simulation data.
+4.3.3 Analysis of the CG dynamics
+Following the same procedure as in the previous examples, we learned a $2\times 2$ diffusion matrix in the CG space, but this time, we tested out a full non-diagonal diffusion field.
+Figure 11 shows the four elements of the learned diffusion matrix.
+In addition, the left panel of Figure 12 depicts the free energy surface learned by the KDE method, which is in satisfactory agreement with the reference one in Figure 2.
+Figure 11: Components of the learned diffusion covariance matrix $a^{\xi}$ for Chignolin in its two-dimensional TICA space (note that the off-diagonal elements are symmetric).
+Figure 12: Free energy surface in the two-dimensional TICA space for Chignolin, as learned by the KDE estimator on the left, and obtained from a histogram of the CG dynamics on the right.
 […]
+As we are mainly interested in kinetic properties, we compute a new gEDMD model on the propagated CG dynamics, and re-compute the associated eigenvalues and eigenvectors.
+The result of a PCCA+-analysis indicates that the correct meta-stable sets are recovered, as shown in the left panel in Figure 13.
+Likewise, the leading implied timescales estimated from the simulated CG dynamics are in good agreement with those of the original gEDMD model $\hat{\mathbf{L}}_{r}$ and the re-scaled MSM timescales, both estimated from the original simulation data, as shown in the right panel of Figure 13.
+Similar to the previous example, we also generate a separate trajectory based on a constant diffusion according to the average of the learned diffusion.
+We find that transition timescales for the constant diffusion are not well fitted to the reference.
+This is likely due to the fact that the variation of the diffusion field is stronger than in the previous example, and taking the mean leads to a slower system.
+This result confirms the need to learn a state-dependent diffusion field in the CG space to achieve kinetic consistency.
+Figure 13: Kinetic consistency of the learned CG model for Chignolin.
+Left: PCCA+ states obtained from simulating the learned CG model.
+Right: slowest finite timescales of the system calculated using an approximation of the generator from the reference dataset (blue) and from the propagated CG dynamics (state-dependent diffusion in orange, constant diffusion in green).
+We also compare to re-scaled timescales from a Markov state model on the original simulation data (black).
 5 Discussion
-We presented the use of kernel-based approximation to the Koopman generator – kgEDMD – to determine meta stable sets and to learn the parameters of CG models for stochastic systems.
-For reversible systems, we presented a single learning pipeline to identify both the potential energy and the diffusion governing the effective SDE in CG space.
-By means of two examples, a two-dimensional model potential and the alanine dipeptide, we showed that the learned effective dynamics is able to reproduce both thermodynamic and kinetic quantities of the full dynamics with high accuracy.
-In the molecular example, we found that the CG dynamics based on the fully state-dependent diffusion performed comparably to a constant-diffusion model using the average value of the learned diffusion.
-It should be noted, however, that determining this average value requires to learn the optimal state-dependent model in the first place.
-However, these observations can serve as a basis for a bottom-up construction of the effective diffusion, starting from a simple constant model, and increasing the complexity incrementally.
-Additional future research topics include improvements to the efficiency of the method, using, for example, sparse approximations using random Fourier features 45 .
+We presented a novel approach to learn kinetically consistent coarse grained models for stochastic dynamics.
+We have introduced a learning method for the effective diffusion field in CG space, and shown how the kinetic properties of the CG dynamics can be evaluated by exploiting models for the Koopman generator (gEDMD algorithm).
+We have also shown that random Fourier features provide an efficient and flexible parametrization for both the effective diffusion and the gEDMD model.
+By means of three examples, a two-dimensional model potential and two datasets of molecular dynamics simulations, we showed that the effective dynamics in low-dimensional reaction coordinate spaces are able to reproduce both thermodynamic and kinetic quantities of the full dynamics accurately.
+For the molecular examples, we have relied on the overdamped assumption to parametrize reversible CG dynamics.
+We have seen that this assumption leads to a uniform acceleration of the CG dynamics compared to the full system.
+The re-scaling factor can be estimated numerically by comparing the gEDMD model to a kinetic model that does not rely on the overdamped assumption.
+We used MSMs in this paper, but note that a more general EDMD model (e.g.
+using random features) would work just as well.
+In this study, we used long equilibrium simulations to train CG models.
+However, one of the appealing aspects of the generator EDMD approach is that it only requires Boltzmann samples.
+As has been pointed out in previous studies, these samples can also be obtained from biased sampling simulations 42, 43 , or by employing generative models 44 .
+Among other topics, future work will focus on applying the formalism to higher-dimensional and more transferrable CG coordinates, for example C-alpha models.
+Another topic is the construction of CG models that can explicitly account for the underdamped structure of the full system, or that can incorporate memory terms, which were entirely disregarded in our study.
 Moreover, one can also try to simultaneously optimize the CG mapping $\xi$ along with the parameters of the CG model, for instance by balancing the VAMP score versus the complexity of the CG model.
+Acknowledgments
+We thank the Theoretical and Computational Biophysics Group at Freie Universität Berlin for sharing the simulation data of the Chignolin mini-protein.
 Data Availability Statement
@@
 $X_{t}$ stochastic process
-$k,\Phi$ kernel and associated feature map
-$\mathbb{H}$ reproducing kernel Hilbert space
 $\mathcal{K}^{\tau}$ Koopman operator with lag time $\tau$
@@
 $f^{\xi}_{\rm lmf}$, $a^{\xi}_{\rm loc}$ local mean force and local diffusion
-$\nabla_{1}f$ gradient of $f$ with respect to the first argument
-$\nabla_{2}f$ gradient of $f$ with respect to the second argument
-$A\cdot_{|a,b}B$ contraction of arguments $a$-th of $A$ and $b$-th of $B$
+$A\cdot_{|i,j}B$ contraction of dimensions $i$ and $j$ of arrays $A$ and $B$.
 6.2 VAMP-score
-We tune hyper-parameters of the proposed method based on VAMP variational principle proposed in 27, stating that for reversible systems, the $k$ dominant eigenvalues of the Koopman generator can be obtained by a minimization problem
+We tune hyper-parameters of the proposed method based on VAMP variational principle proposed in 26 , stating that for reversible systems, the $k$ dominant eigenvalues of the Koopman generator can be obtained by a minimization problem
 $\sum_{i=1}^{k}\lambda_{i}=\min_{\phi_{0},...,\phi_{k}}\sum_{i=1}^{k}\left\langle\phi_{i},\,\mathcal{L}\phi_{i}\right\rangle_{\mu}$ (36)
 where the $\phi_{i}$ are orthogonal functions.
-We use this variational principle to optimize the kernel bandwidth in the context of our proposed method.
+We use this variational principle to optimize the kernel bandwidth giving rise to the spectral measure used in the context of our proposed method.
 To do this robustly and avoid overfitting, we make use of standard cross validation scheme by introducing $40\%$ of dataset as test set.
 Figure 14 shows the result for the lemon slice example, using Gaussian and periodic Gaussian kernels.
-Figure 9: VAMP-score analysis for the Lemon slice example using periodic Gaussian kernel on the left and Gaussian kernel on the right
-We applied the same procedure for optimizing the bandwidth for the example of alanine dipeptide.
-The only difference is that due to the relative complexity of this example compared to the Lemon slice, and requiring to incorporate more data for the optimization, we use random Fourier features (RFF) to reduce the size of basis functions, leading to a low-rank approximation of the generator.
-Figure 10 shows the result for optimization of the bandwidth.
-Figure 10: VAMP-score analysis for alanine dipeptide.
+Figure 14: VAMP-score analysis for the Lemon slice example using periodic Gaussian kernel on the left and Gaussian kernel on the right.
+We applied the same procedure for optimizing the bandwidth for alanine dipeptide and Chignolin. Figure 15 shows the result for optimization of the bandwidth.
+Figure 15: VAMP-score analysis for alanine dipeptide on the left and Chignolin on the right.
+As the figure indicates, there exist a range of the bandwidth $\gamma$ that we can safely choose the bandwidth from.
 6.3 Simulation Settings for Alanine Dipeptide